### v1

Abstract
Given an Abelian group $\mathcal{G}$, a Boolean-valued function $f:\mathcal{G}\to\{-1,+1\}$, is said to be $s$-sparse, if it has at most $s$-many non-zero Fourier coefficients over the domain $\mathcal{G}$. In a seminal paper, Gopalan et al. [15] proved “Granularity” for Fourier coefficients of Boolean valued functions over $\mathbb{Z}_{2}^{n}$, that have found many diverse applications in theoretical computer science and combinatorics. They also studied structural results for Boolean functions over $\mathbb{Z}_{2}^{n}$ which are approximately Fourier-sparse. In this work, we obtain structural results for approximately Fourier-sparse Boolean valued functions over Abelian groups $\mathcal{G}$ of the form, $\mathcal{G}:=\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$, for distinct primes $p_{i}$. We also obtain a lower bound of the form $1/(m^{2}s)^{\lceil\varphi(m)/2\rceil}$, on the absolute value of the smallest non-zero Fourier coefficient of an $s$-sparse function, where $m=p_{1}\cdots p_{t}$, and $\varphi(m)=(p_{1}-1)\cdots(p_{t}-1)$. We carefully apply probabilistic techniques from [15], to obtain our structural results, and use some non-trivial results from algebraic number theory to get the lower bound.
We construct a family of at most $s$-sparse Boolean functions over $\mathbb{Z}_{p}^{n}$, where $p>2$, for arbitrarily large enough s, where the minimum non-zero Fourier coefficient is $o(1/s)$. The “Granularity” result of Gopalan et al. implies that the absolute values of non-zero Fourier coefficients of any $s$-sparse Boolean valued function over $\mathbb{Z}_{2}^{n}$ are $\Omega\left({1}/{s}\right)$. So, our result shows that one cannot expect such a lower bound for general Abelian groups.
Using our new structural results on the Fourier coefficients of sparse functions, we design an efficient sparsity testing algorithm for Boolean function, which tests whether the given function is $s$-sparse, or $\epsilon$-far from any sparse Boolean function, and it requires $\mathsf{poly}((ms)^{\varphi(m)},1/\epsilon)$-many queries. Further, we generalize the notion of degree of a Boolean function over the Abelian group $\mathcal{G}$. We use it to prove an $\Omega(\sqrt{s})$ lower bound on the query complexity of any adaptive sparsity testing algorithm.
keywords:
Fourier coefficients, sparse, Abelian, granularity
1 Introduction
Boolean functions are fundamental objects of study in computer science. For a discrete domain $\mathcal{D}$, a Boolean function $f:\mathcal{D}\to\{+1,-1\}$ models a decision task where each member of $\mathcal{D}$ is classified into one of two classes. Boolean functions play a vital role in the study of digital circuits and computer hardware. They are also significant in the study of algorithms and complexity, particularly in problems where the set $\mathcal{D}$ of instances is endowed with an algebraic structure. Examples of such problems include matrix multiplication and polynomial evaluation.
The case of Boolean function complexity with $\mathcal{D}=\mathbb{Z}_{2}^{n}$ has been widely studied. These functions are often analyzed in connection with their Fourier transform (see the archived version of the paper, cite) and a significant amount of research has focused on the structural properties of Fourier spectra of important classes of these functions. One such class that this work focuses on is that of Fourier-sparse functions. These are functions with only a few non-zero Fourier coefficients, formally defined in the archived version of the paper, cite. We will denote by $s_{f}$ the Fourier sparsity of a Boolean function $f$. Fourier sparsity and Fourier-sparse functions have known connections with a variety of areas of Boolean function analysis and computational complexity like property testing [15], learning theory [21, 3], distance estimation [38] and communication complexity [25, 34, 20, 29, 24]. These connections provide enough motivation to comprehend the structure of Fourier coefficients for Fourier-sparse Boolean functions.
In this work, we extend the study of Fourier-sparse Boolean functions to the domains $\mathcal{D}$, which are Abelian groups of the form $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$, where $p_{1},\ldots,p_{t}$ are distinct prime numbers. Boolean functions over general Abelian groups have been studied in both mathematics and computer science. A celebrated result regarding such functions is Chang’s Lemma [10]. Chang’s lemma over $\mathbb{Z}_{2}^{n}$ has found numerous applications in complexity theory and algorithms [6, 9], analysis of Boolean functions [17, 34], communication complexity [34, 22], extremal combinatorics [13], and many more. Recently, [8] improved Chang’s lemma over $\mathbb{Z}_{2}^{n}$ for some special settings of parameters, where Fourier sparsity played a crucial role. One motivation to study Fourier sparsity over a broader class of Abelian groups is to investigate possible generalizations of their bounds to those groups.
Fourier analysis over finite Abelian groups in cryptography. For the past three decades, the field of cryptography has been utilizing concepts derived from Fourier analysis, specifically over finite Abelian groups. Akavia, Goldwasser, and Safra [2] have combined some of these concepts to develop a comprehensive algorithm that can detect “large” Fourier coefficients of any concentrated function on finite Abelian groups, and compute a sparse approximation for the same. This algorithm has gained significant attention within the cryptography community, especially regarding the notion of “bit security” of the discrete logarithm problem (DLP), RSA, and learning with errors (LWE) problems; see [27, 14, 1]. In particular, the “nice” structural results on the Fourier coefficients of a Boolean-valued function over the general Abelian group are of utmost importance and interest from a crypto-theoretic point of view.
Interestingly, there are strong relationships between learning, sparsity, and sampling, in the context of Fourier-sparse Boolean functions. They have been rigorously studied in [31, 32, 33]. In [31], the authors asked the following:
Question 1.1.
What can be said about the structure of the Fourier coefficients of a Boolean function $f$ over an Abelian group $\mathcal{G}=\mathbb{Z}_{p}^{n}$, where the support is significantly smaller compared to $\mathcal{G}$?
Gopalan et al. [15] proved that any non-zero Fourier coefficient of a Boolean function over $\mathbb{Z}_{2}^{n}$ with Fourier sparsity at most $s_{f}$, is at least $\frac{1}{s_{f}}$ in its absolute value. This gave a satisfactory answer of 1.1 over $\mathbb{Z}_{2}^{n}$. Furthermore, they proved robust versions of their result for functions which are approximately Fourier-sparse. Finally, their structural results were used to design a sample-efficient algorithm to test whether a function is Fourier-sparse.
In our work, we undertake the same task for Boolean functions over Abelian groups of the form $\mathbb{Z}_{p}^{n}$, for a prime $p\geq 3$. We prove lower bounds on the absolute value of any non-zero coefficient in terms of $s_{f}$ and $p$. We also prove a tightness result that complements our lower bounds. In particular, our bound implies that a lower bound of $\frac{1}{\Theta(s_{f})}$ that [15] showed does not hold anymore for $\mathbb{Z}_{p}^{n}$. Finally, we use our bounds to design a testing algorithm for Fourier-sparse Boolean functions over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$. This is probably the first time, we have advanced on understanding some structure on the Fourier-sparse coefficients over a general Abelian group, and shed some light on 1.1.
It is well-known that any Abelian group is isomorphic to a group of the form $\mathbb{Z}_{p_{1}^{n_{1}}}\times\cdots\times\mathbb{Z}_{p_{t}^{n_{t}}}$ where $p_{1},\ldots,p_{t}$ are prime numbers and $n_{1},\ldots,n_{t}$ are positive integers. Unfortunately, our techniques fall short of handling even a simpler case of $\mathbb{Z}_{p^{2}}$, fundamentally because of the absence of a linear structure on the solution space of systems of equations over $\mathbb{Z}_{p^{2}}$. Therefore, we are leaving the task of investigating the existence of similar bounds and algorithmic results when the domain of the function is a general Abelian group to future research.
Why care about Fourier-sparse Boolean functions over Abelian groups?
There has been a considerable amount of interest in studying the complexity of reconstructing or learning functions of the form $f:\mathcal{D}\to\mathbb{C}$, where $\mathcal{D}$ is a known domain (more general than a hypercube, such as a general finite Abelian group), and $f$ is Fourier-sparse; see [33, 28, 26, 11, 37]. Fourier-sparse functions over various finite Abelian groups have gained much interest with the advancement in sparse Fourier transform algorithms [18, 19, 5]. These algorithms improve the efficiency of the standard Fast Fourier Transform algorithms by taking advantage of the sparsity itself. To reliably use sparse Fourier transform algorithms, it is beneficial to have a way to test if a function is $s$-sparse or, more generally, to estimate the distance of a function to the closest $s$-sparse function. In this work, we consider the problem of non-tolerant sparsity-testing of Boolean functions over finite Abelian groups.
Finally, apart from mathematical curiosity and potential cryptographic applications (as mentioned previously), structural results on Fourier-sparse functions $f:\mathbb{Z}_{N}\to\mathbb{C}$, for some $N\in\mathbb{N}$, have also found algorithmic applications in SOS-optimization and control theory. These applications have further implications in certifying maximum satisfiability (MAX-SAT) and maximum k-colorable subgraph (MkCS) problems; see [12, 36, 35].
1.1 Our results
Throughout the article, we will be working with the Abelian group $\mathcal{G}:=\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ where $p_{i}$ are primes. Let $f:\mathcal{G}\to\{-1,+1\}$, and $f(x)=\sum_{\chi\in\hat{\mathcal{G}}}\widehat{f}(\chi)\chi(x)$ be the Fourier transform of $f$, where $\hat{\mathcal{G}}$ is the set of characters of the Abelian group $\mathcal{G}$.
We say a Boolean function is $s$-sparse, if it has at most $s$ non-zero Fourier coefficients. [15] proved that for any $s$-sparse Boolean functions over $\mathbb{Z}_{2}^{n}$, the magnitude of the Fourier coefficients are $k$-granular where $k=\lceil\log_{2}s\rceil+1$. A real number is $k$-granular if it is an integer multiple of $1/2^{k}$. One wonders whether such a phenomenon still holds over a more general group $\mathcal{G}$.
This notion of granularity made sense over $\mathbb{Z}_{2}^{n}$, since in this case, all the Fourier coefficients are rational (and hence real) numbers. But when the domain of the function is a general group $\mathcal{G}$, the Fourier coefficients are necessarily complex numbers. So, we would like to suitably define granularity, and show that such a property still holds for $s$-sparse Boolean-valued functions over $\mathcal{G}$. Our first conceptual contribution in this paper is to generalize the notion of granularity appropriately.
Definition 1.1.
(Granularity) A complex number is said to be $k$-granular or has granularity $k$ with respect to $\mathbb{Z}_{p}$ if it is of the form $\frac{g(\omega_{p})}{p^{k}}$, where $g(X)\in\mathbb{Z}[X]$ and $\omega_{p}$ is a primitive $p^{th}$ root of unity.
More generally, a complex number is said to be $(m_{1},\ldots,m_{t})$-granular with respect to $(\mathbb{Z}_{p_{1}},\dots,\mathbb{Z}_{p_{t}})$ if it is of the form $\frac{g(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}$, where $g(X_{1},\ldots,X_{t})\in\mathbb{Z}[X_{1},\ldots,X_{t}]$ and $\omega_{p_{i}}$ is a primitive $p_{i}^{th}$ root of unity, $i\in[k]$.
Note that, this goes well with the definition of granularity in [15] for the case of $\mathbb{Z}_{2}$ as $\omega_{2}$ is either $+1$ or $-1$ and hence $g(\omega_{2})$ is an integer for any $g(X)\in\mathbb{Z}[X]$.
A function $f:\mathcal{G}\to\{-1,+1\}$ is said to be $(m_{1},\dots,m_{t})$-granular with respect to $(\mathbb{Z}_{p_{1}},\ldots,\mathbb{Z}_{p_{t}})$ if each Fourier coefficient of $f$ is $(m_{1},\dots,m_{t})$-granular.
We will also need a robust version of the definition of granularity of a complex number.
Definition 1.2.
($\mu$-close to $k$-granular) A complex number $v$ is said to be $\mu$-close to $k$-granular with respect to $\mathbb{Z}_{p}$ if $|v-\frac{g(\omega_{p})}{p^{k}}|\leq\mu$, for some non-zero polynomial $g(X)\in\mathbb{Z}[X]$.
Note that a similar notion can be defined for the case of $\mu$-close to $(m_{1},\ldots,m_{t})$-granular with respect to $(\mathbb{Z}_{p_{1}},\dots,\mathbb{Z}_{p_{t}})$.
Now, we are ready to formally state our two main structural results. All our results hold for Boolean-valued functions over the more general Abelian group $\mathcal{G}$. But for simplicity of presentation and ease of understanding, we first present the result for Boolean-valued functions over $\mathbb{Z}_{p}^{n}$. Later, in Appendix F, we present more general versions of our results.
Our first theorem says that for any Boolean-valued function over $\mathbb{Z}_{p}^{n}$ that is close to being sparse, all its large Fourier coefficients are close to being granular. This is a generalization of the structural theorem of [15, Theorem 3.3], which was proved over $\mathbb{Z}_{2}^{n}$.
Theorem 1.3 (Structure theorem 1).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean-valued function and let $B$ be the set of characters corresponding to the set of $s$-largest Fourier coefficients of $f$ (in terms of magnitude). If $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$ then for all $\chi\in B$, $\widehat{f}(\chi)$ is $\frac{\mu}{\sqrt{s}}$-close to $\lceil\log_{p}s\rceil+1$-granular.
For a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$ (where $B$ is the set of characters corresponding to the set of $s$ largest coefficients of $f$) can also be stated as “there is an $s$-sparse function $g:\mathbb{Z}_{p}^{n}\to\mathbb{C}$ with the $\ell_{2}$-distance between $f$ and $g$ is at most $\sqrt{\mu}$”. But note that this does not guarantee that there is an $s$-sparse Boolean-valued function $g:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ with $\ell_{2}$-distance between $f$ and $g$ being at most $\sqrt{\mu}$. However, our second theorem proves that one can indeed find an $s$-sparse Boolean-valued function in a close enough vicinity, thus generalizing [15, Theorem 3.4].
Theorem 1.4 (Structure theorem 2).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean-valued function and let $B$ be the set of characters corresponding to the set of $s$-largest Fourier coefficients of $f$ (in terms of magnitude). If $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$, with $\mu\leq\frac{1}{8(p^{2}s)^{p-1}}$ then there exists an $s$-sparse Boolean-valued function $F:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ with $\ell_{2}$ distance between $f$ and $F$ is at most $\sqrt{2}\mu$.
Theorem 1.3 and Theorem 1.4 can be suitably generalized to functions from $\mathcal{G}$ to $\{-1,+1\}$. But for the clarity of presentation, we state the generalized theorems and present their proofs in Appendix F.
One important corollary of Theorem 1.3 is that for any $s$-sparse Boolean function $f:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$ the non-zero Fourier coefficients has magnitude at least $1/2^{k}$, where $k=\lceil\log_{2}s\rceil+1$. Unfortunately, such a simple corollary cannot be claimed for $s$-sparse functions $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ if $p\neq 2$. The main reason is that the definition of granularity is for complex numbers, rather than real numbers and hence such a lower bound cannot be directly deduced. However, borrowing results from algebraic number theory, we can obtain a lower bound on the Fourier coefficients of Boolean-valued functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$ for any arbitrary prime $p$.
Theorem 1.5 (Fourier coefficient lower bound).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, with Fourier sparsity $s_{f}$. Then, for any $\chi\in\operatorname{supp}(f)$, we have $|\hat{f}(\chi)|\geq\frac{1}{(p^{2}s_{f})^{\lceil(p-1)/2\rceil}}$
{remark*}
One can also prove a weaker lower bound of the form $1/((s_{f}+1)\sqrt{s_{f}})^{s_{f}}$, which is $p$-independent; for details, see Theorem C.8. Observe that the lower bound in Theorem 1.5 is much lower than $1/s_{f}$. One may wonder how tight our result is. It is known that $1/s$ is tight for the case when $p=2$. For example, consider the function $AND:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$. Its non-empty Fourier coefficients are either $\frac{1}{2^{n-1}}$ or $-\frac{1}{2^{n-1}}$, while the empty (constant) coefficient being $1-\frac{1}{2^{n-1}}$.
To our pleasant surprise, we construct $s$-sparse Boolean-valued functions over $\mathbb{Z}_{p}^{n}$, for $p\geq 5$, such that they have nonzero Fourier coefficients with absolute value being $o(1/s)$.
Theorem 1.6 (Small Fourier coefficients).
For every prime $p\geq 5$, and large enough $n$, there exist a positive constant $\alpha_{p}$ that depends only on $p$ and a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ with Fourier sparsity $s_{f}$ satisfying the following property:
$\min_{\chi\in\operatorname{supp}(f)}\left|\widehat{f}(\chi)\right|\leq 1/s_{f}^{1+\alpha_{p}}.$
We prove a generalized version of the lower bound result (Theorem 1.5) for Boolean-valued functions over $\mathcal{G}$. The above example (from Theorem 1.6) can also be easily extended to obtain Boolean-value functions over $\mathcal{G}$ demonstrating similar bounds on the absolute value of the non-trivial Fourier coefficients.
Finally, we design efficient algorithms for testing whether a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is $s$-sparse or “far” from $s$-sparse-Boolean. To state our results we need to define what we mean by a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse.
Definition 1.7.
A function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse-Boolean if for every $s$-sparse function $g:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ the $\ell_{2}$-distance of $f$ and $g$ is at least $\sqrt{\epsilon}$. [^1] In property testing usually the distance measure used is Hamming distance between two Boolean functions. But since we are using $\ell_{2}$ distance in our other theorem, so for ease of presentation, we have defined the farness in terms of $\ell_{2}$ instead of Hamming distance. Also, note that for a pair of Boolean-valued functions the square of the $\ell_{2}$ distance and Hamming distance are the same up to a multiplicative factor of $4$, see Lemma A.15.
We say that an algorithm (property tester) $\mathcal{A}$  $\epsilon$-tests $\mathcal{C}$, for a class of functions $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, if given access to the truth table of a function $f$, whether $f\in\mathcal{C}$, or $f$, is “$\epsilon$-far from $\mathcal{C}$” can be tested using $\mathcal{A}$ with success probability (called the confidence) $\geq 2/3$. The number of queries to the truth-table of $f$ made by $\mathcal{A}$ is called the query complexity of $\mathcal{A}$.
Using the structure theorems (Theorem 1.3 and Theorem 1.4), we prove the following theorem which tests sparsity of a Boolean-valued function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$.
Theorem 1.8 (Testing $s$-sparsity).
For a fixed prime $p$, there is a non-adaptive $\mathsf{poly}(s,1/\epsilon)$ query algorithm with confidence $2/3$, which tests whether a given function $f:\mathbb{Z}_{p}^{n}\rightarrow\{-1,+1\}$, is $s$-sparse or $\epsilon$-far from $s$-sparse-Boolean.
We can also obtain a similar testing algorithm for sparsity, for a Boolean-valued function over $\mathcal{G}$. The generalized algorithm is discussed in Appendix G.
We complement our result by showing a query-complexity lower bound for sparsity-testing algorithms. Gopalan et al. [15] gave a $\Omega(\sqrt{s})$ lower bound for $s$-sparsity testing algorithms over $\mathbb{Z}_{2}^{n}$. An important component of their proof was to cleverly use an alternative notion of degree (borrowed from [7]) of a Boolean function over $\mathbb{Z}_{2}$. We also give a similar lower bound over $\mathbb{Z}_{p}^{n}$, by appropriately generalizing the useful notion of degree. For details on the definition of degree, see proof idea of Theorem 1.9 in Section 1.2 and Appendix H.
Theorem 1.9.
For Boolean valued functions on $\mathbb{Z}_{p}^{n}$, to adaptively test $s$-sparsity, the query lower bound of any algorithm is $\Omega(\sqrt{s})$.
Theorem 1.9 can be generalized for Boolean valued functions on $\mathcal{G}$, which will give us the same lower bound.
1.2 Proof ideas
In this section, we briefly outline the proof ideas of our main theorems. Although some of the proofs are indeed inspired by [15] (which worked over $\mathbb{Z}_{2}$), the proof techniques does not directly generalize over $\mathbb{Z}_{p}$. So, we will first discuss the proof ideas, and then clarify the differences between [15] and our techniques (see Section 1.2.1). While doing so, we will try to convey the hurdles for generalizing over $\mathbb{Z}_{p}$. Let us first sketch the proof of Theorem 1.3.
Proof idea of Theorem 1.3. Our goal is to show that if $f$ is $\mu$-close to some $s$-sparse complex-valued function in $\ell_{2}$, then there exists a non-zero polynomial $g(X)\in\mathbb{Z}[X]$ such that the following properties hold.
  1. 1.
    The sum of the absolute values of its coefficients is at most $p^{k}$, where $k:=\lceil\log_{p}s\rceil+1$.
  2. 2.
    The distance between the absolute value of each non-zero Fourier coefficient of $f$ and $|g(\omega_{p})|/p^{k}$ is at most $\mu/\sqrt{s}$.
To show the above, we first utilize a probabilistic method to prove that for each character $\chi_{i}\in B$ in the Fourier support of $f$, there exists a matrix $A\in\mathbb{Z}_{p}^{k\times n}$, and a column vector $b\in\mathbb{Z}_{p}^{n\times 1}$, such that – (1) $\chi_{i}$ is a solution of the linear system $A\chi=b$, and (2) no other character in the Fourier support of $f$ is a solution of $A\chi=b$, where $B$ be the set of $s$-largest Fourier coefficients of $f$. After establishing the existence of such $A$ and $b$, we consider the Fourier transform of the projection operator for the solution space of $A\chi=b$ (see Definition B.1 and Lemma B.2). The projection operator, as the name suggests, is an operator that projects $\mathbb{Z}_{p}^{n}$ onto a linear subspace which yields a partition of the Fourier spectrum of $f$. Then we show that the $\ell_{2}$ Fourier weight of $S\cap H$, i.e., $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$ is upper bounded by $\mu/\sqrt{s}$, where $S=\overline{B}$, and $H$ is a coset of $A^{\perp}$ that are solutions to the system of linear equations $A\chi=b$. For details, see Appendix B.
[…]
As stated earlier, we use a non-trivial result from algebraic number theory (Theorem 2.1), which states that if $f\in\mathbb{Z}[x]$, such that $f(\omega_{n})\neq 0$, where $\omega_{n}$ be a primitive root of unity, then, $|\prod_{i\in\mathbb{Z}_{n}^{\ast}}f(\omega_{n}^{i})|\geq 1$. We also use the fact that the sum of the absolute values of the coefficients of $g$ is at most $p^{k}$, to get an upper bound on the quantities $|g(\omega_{p}^{i})|$, for any $i\in[p-1]$. Combining these two facts, we obtain our lower bound; for details see Section B.
Proof idea of Theorem 1.4. We first show that the given function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, that is $\mu$-close to some $s$-sparse complex-valued function in $\ell_{2}$, can be written as the sum of two functions $F$ and $G$, where the Fourier coefficients of $F$ are $\lceil\log_{p}s\rceil+1$-granular and the absolute value of the Fourier coefficients of $G$ are upper bounded by $\mu/\sqrt{s}$, where $p$ is an odd prime. This follows from Theorem 1.3. Then we show that the range of $F$ is $\{-1,+1\}$, which uses the following facts:
  1. 1.
    $(F+G)^{2}=f^{2}=1$ and
  2. 2.
    $F^{2}$ is $12\lceil\log_{p}s\rceil+$ 1-granular.
We compute $\mathbb{E}[G(x)^{2}]$ in order to find an upper bound on the Fourier coefficients of $H:=G(2f-G)$, which helps us to conclude that $\widehat{F^{2}}(\chi)=0$ for all $\chi\neq\chi_{0}$, and $\widehat{F^{2}}(\chi_{0})=1$, where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathbb{Z}_{p}^{n}$. Then we complete the proof by showing that the $\Pr_{x}[x\in\mathbb{Z}_{p}^{n}|f(x)\neq g(x)]$ is $\leq\mu^{2}/2$, which implies that $F$ is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$ by Lemma A.15. This idea has also been employed in  [15]. We have also extended this proof to a more general Abelian group $\mathcal{G}$, (see Appendix F).
Proof idea of Theorem 1.6. The example that we construct to prove Theorem 1.6 is a simple one. The crucial observation in this regard is that there are functions from $\mathbb{Z}_{p}$ to $\{-1,+1\}$, whose Fourier coefficients are smaller than $1/p$. In this work, we work with one such function $\mathbb{I}_{\geq\frac{p+1}{2}}:\mathbb{Z}_{p}\to\{-1,+1\}$. We claim that the composition function $AND_{n}\circ\mathbb{I}_{\geq\frac{p+1}{2}}$ is one such desired function from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$, such that its minimum Fourier coefficient is less than $1/p^{n}$. Here, we assume a trivial bound of $p^{n}$ on the Fourier sparsity of the function $AND_{n}\circ\mathbb{I}_{\geq\frac{p+1}{2}}$.
Proof idea of Theorem 1.8. We sketch the algorithmic idea over $\mathbb{Z}_{p}^{n}$, where $p$ is an odd prime; this idea can be canonically extended to more general Abelian groups $\mathcal{G}$. The main idea of the algorithm is to partition the set of characters into buckets and estimate the weights of the individual buckets (that is the sum of the squares of the absolute values of the Fourier coefficients corresponding to the characters in the buckets). We know from Theorem 1.5 that all the coefficients of an $s$-sparse function are at least as large as $1/(p^{2}s)^{\lceil(p-1)/2\rceil}$. So, we are certain that if the weights of the buckets can be approximated within an additive error or $\tau/3$, where $\tau\geq 1/(p^{2}s)^{p-1}$, then in the case the function is $s$-sparse, not more than $s$ of the buckets can have weight more than $\tau/2$. On the other hand, we will show that if the function $f$ is $\epsilon$-far from any $s$-sparse Boolean function then with a high probability at least $(s+1)$ buckets will have weight more than $\tau$, making the estimated weight at least $32\tau/$ 3; see Lemma E.5.
The challenge is that estimating the weights of the buckets is not easy if the characters the randomly partitioned into buckets. Here we use the ideas from Gopalan et al [15] and appropriately modify them to handle the technicalities of working with $\mathbb{Z}^{n}_{p}$. We choose the buckets carefully. The buckets corresponds to the cosets of $H^{\perp}$ in $\mathbb{Z}_{p}^{n}$, where $H$ is a random subspace of $\mathbb{Z}_{p}^{n}$ of dimension, $t=\Theta(s^{2})$. For such kinds of buckets, we show that estimation of the weight can be done using a small number of samples. We also need to use the concept of ‘random shift’, see Definition E.12, to avoid the corner case of characters being put into the trivial bucket.
Unlike [15], it becomes a bit more challenging to prove that if $f$ is $\epsilon$-far from any $s$-sparse Boolean function over $\mathbb{Z}_{p}^{n}$, then with high probability at least $(s+1)$ buckets will have weight more than $\tau$. Since we partition the set of characters by cosets, the events whether two characters land in the same bucket (that is same coset) are not independent – since two characters (in the case $p\neq 2$) can be scalar multiple of each other; this is where some additional care is required (which was not the case in [15]). Under random shifts and case-by-case analysis, we can show that the two events are not correlated, i.e., the covariance of the corresponding indicator variables of the two events is 00; see Lemma E.13. Thus, we can use Chebyshev’s inequality and then Markov’s inequality to bound the number of buckets that can be of low weight, or in other words prove that the number of “heavy” weight buckets is more than $s+1$. Due to lack of space, the proof of Theorem 1.8 is given in Appendix E.
Proof idea of Theorem 1.9. In [15], Gopalan et al. proved a query lower bound over $\mathbb{Z}_{2}^{n}$, by using a natural notion of degree of a Booelan function, denoted $\deg_{2}$. They crucially used the fact that for a Boolean function $f$ over $\mathbb{Z}_{2}$, $2^{\dim(f)}\geq s_{f}\geq 2^{\deg_{2}(f)}$; this was originally proved in [7]. To define the degree, let us consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a subspace of $\mathbb{Z}_{p}^{n}$ as defined as
$V_{b,r_{1},\ldots,r_{k}}\;:=\;\{x\in\mathbb{Z}_{p}^{n}:r_{j}\cdot x\,=\,b_{j}\leavevmode\nobreak\ \pmod{p}\ \forall j\in[t]\}\;.$
Then the degree over $\mathbb{Z}_{p}$ of $f$, denoted by $\deg_{p}$, is defined in the following way.
$\deg_{p}(f)\;=\;\max_{\ell}\{\ell\;=\dim(V_{b,r_{1},\ldots,r_{t}})\;:\;s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}\;=\;p^{\dim(V_{b,r_{1},\ldots,r_{t}})}\},$
where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$. [15] defined the degree over $\mathbb{Z}_{2}^{n}$ via similar restrictions. However, [15] argued that this is a natural definition of the degree over $\mathbb{Z}_{2}^{n}$. To argue, observe that $\widehat{f}=\frac{1}{2^{n}}H_{n}f$, where $H_{n}$ is the $2^{n}\times 2^{n}$ Hadamard matrix, when $x_{i}\in\mathbb{Z}_{2}^{n}$ are seen in lexicographic order. Consider the restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ that takes the value $1$ at those points such that each entry of $\widehat{f}|_{V_{b,r_{1},\ldots,r_{t}}}$ is nonzero. Then $\deg_{2}$ can be defined as the dimension of the subspace $V_{b,r_{1},\ldots,r_{t}}$ which is largest amongst them! In that case, all the Fourier coefficients of $f|_{V_{b,r_{1},\ldots,r_{t}}}$ are nonzero.
Interestingly, we can also show that the above definition of $\deg_{p}$ is natural, mainly because $\widehat{f}=\frac{1}{p^{n}}V_{n}f$, where $V_{1}$ is a $p\times p$ Vandermonde matrix $V_{1}$, whose $(i,j)$-th entry is $(V_{1})_{i,j}:=\omega_{p}^{(i-1)(j-1)}$, and $V_{n}$ is defined by taking $n$ many Kronecker products of $V_{1}$, i.e., $V_{n}:=V_{1}^{\otimes n}=V_{1}\otimes\cdots\otimes V_{1}$. Note that $H_{n}=V_{n}$, over $\mathbb{Z}_{2}^{n}$. Similar to [7], one can also show that $p^{\dim(f)}\geq s_{f}\geq p^{\deg_{p}(f)}$ (Lemma I.3). This plays a crucial role in the proof.
We first define two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ on the set of Boolean valued functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$. Let us choose a random $t$-dimensional subspace $H$ of $\mathbb{Z}_{p}^{Ct}$, for some parameter $C$ (to be fixed later). Let $\mathcal{C}$ be the set of all cosets of $H$. There are 2 main steps as follows.
  1. 1.
    We construct random functions $f$ by making $f$ a constant on each coset of $\mathcal{C}$. The constant is chosen randomly from $\{-1,+1\}$. We call this probability distribution $\mathcal{D}_{yes}$.
  2. 2.
    We choose a random function $f$ randomly from $\mathbb{Z}_{p}^{Ct}$, conditioned on the fact that $f$ is $2-\tau$ far in $\ell_{2}$ from any function which has $\deg_{p}=t$, where $\tau$ is as defined in Theorem 1.8. Let us call this distribution $\mathcal{D}_{No}$.
We show that if an adaptive query algorithm makes less than $q<\Omega(p^{t/2})$ queries, then the total variation distance $||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}$ between the two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ is $\leq\frac{1}{3}$. This proves that any adaptive query algorithm which distinguishes between $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$, i.e., where $||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}>\frac{1}{3}$, must make at least $\Omega(p^{t/2})$ queries. This essentially proves Theorem 1.9.
1.2.1 Comparison with Gopalan et al. [15]
In this section, we discuss the main differences and points between our proof and the one by Gopalan et al. [15]. Our proofs (and the analysis of the testing algorithm) are more elaborate, subtle, case-dependent, and complicated than [15].
Linear dependence is bad. [15] used the similar idea in Theorem F.2. However, their proof took advantage of the fact that any two distinct non-zero vectors in $\mathbb{Z}_{2}^{n}$ are linearly independent. This is not true for $\mathbb{Z}_{p}^{n}$, as distinct vectors can be scalar multiples of each other; e.g., $(1,\cdots,1)$ and $(2,\cdots,2)$ in $\mathbb{Z}_{p}^{n}$, for any $p\geq 3$. Thus, our proof technique is capable of handling and overcoming these difficulties effectively. This is also why our analysis of the structure theorems and algorithms are very much case-dependent.
Different granularity. In case of $\mathbb{Z}_{2}^{n}$,  [15] defined a function $f:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$ to be $k$-granular, where $k$ is a positive integer, if all of its nonzero Fourier coefficients are an integer multiple of $\frac{1}{2^{k}}$. In the case of $\mathbb{Z}_{p}^{n}$, since the Fourier coefficients are truly complex numbers, we define a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ to be $k$-granular if all of its nonzero Fourier coefficients are of the form $\frac{g(\omega_{p})}{p^{k}}$ (see Definition 1.1), where $g\in\mathbb{Z}[X]$ and $\omega_{p}$ is a primitive $p^{th}$ root of unity.
Algebraic integer lower bounds are harder. While proving a lower bound on the absolute value of the Fourier coefficients, in the case of $\mathbb{Z}_{2}^{n}$, proving a lower bound on the magnitude of a Fourier coefficient  [15] translates to proving a lower bound on $|g(-1)|/2^{k}$ (since $\omega_{2}=-1$). Note that $|g(-1)|\geq 1$, because $g(-1)\in\mathbb{Z}\setminus\{0\}$; hence the lower bound follows. However, for a general prime $p$, $\omega_{p}$ is truly a complex number, and hence, $g(\omega_{p})$ is no longer an integer (and therefore, there is no way to conclude that $|g(\omega_{p})|\geq 1$). We use a non-trivial result from algebraic number theory (Theorem 2.1) to tackle this problem, in the proof of Theorem 1.5. We also crucially use the fact that the sum of the absolute values of the coefficients of $g$ is at most $p^{k}$, to get an upper bound on the quantities $|g(\omega_{p}^{i})|$, for any $i\in[p-1]$. Clearly, this part requires much more non-triviality than [15].
Do not expect a tight linear lower bound. For $p=2$, it follows from  [15] that the lower bound on the absolute value of the Fourier coefficients is equal to $\frac{1}{s_{f}}$. We show in Theorem 1.6 that there exists a family of Boolean valued functions on $\mathbb{Z}_{p}^{n},\ p>3$ such that $\min_{\chi\in\operatorname{supp}(f)}\left|\widehat{f}(\chi)\right|\leq 1/s_{f}^{1+\alpha_{p}}$, hence proving that the lower bound on the absolute value of the Fourier coefficients is not linear in $\frac{1}{s_{f}}$ in case of $p>3$.
Generalizing the notion of degree. For functions from $\mathbb{Z}_{2}^{n}$ to $\{-1,+1\}$, it is known that $\widehat{f}=\frac{1}{p^{n}}H_{n}f$, where $H_{n}$ denotes the Hadamard matrices (see [7]). Then, the degree $\deg_{2}$ is defined naturally in the case of $\mathbb{Z}_{2}^{n}$. The definition of $\deg_{p}$ becomes slightly challenging, since the relation $\widehat{f}=\frac{1}{p^{n}}H_{n}f$ is simply false over $\mathbb{Z}_{p}^{n}$. To define $\deg_{p}$ for Boolean valued functions on $\mathbb{Z}_{p}^{n}$ (see Definition H.6), where $p$ is an odd prime, we have defined the matrix $V_{1}$ using Vandermonde matrix, and have inductively defined the matrices $V_{n}$, where $n$ is a positive integer, via Kronecker product. Then we have shown that $\widehat{f}=\frac{1}{p^{n}}V_{n}f$, which helps us to define $\deg_{p}(f)$. We claim that this is an appropriate generalization of degree, since $H_{n}=V_{n}$, over $\mathbb{Z}_{2}^{n}$.
Finally, we remark that because of the absence of a linear structure on the solution space of systems of equations over arbitrary Abelian groups (e.g. $\mathbb{Z}_{p^{2}}$), we could not generalize this to arbitrary Abelian groups.
2 Lower bound on the Fourier coefficients
In this section, we will prove Theorem 1.5 assuming Theorem 1.3, which gives us a lower bound on the Fourier coefficients of functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$, where $p$ is a prime number. This is a generalization on the granularity of a function $f:\mathbb{Z}_{2}^{n}\to\mathbb{R}$ when the domain of the function is $\mathbb{Z}_{p}^{n}$. We will use the following theorem; for more details see Section A.3.
Theorem 2.1.
For $n\in\mathbb{Z}$, let $\omega_{n}$ be a primitive root of unity. Let $f\in\mathbb{Z}[x]$, such that $f(\omega_{n})\neq 0$. Then,
[…]
Proof 2.2 (Proof of Theorem 1.5).
If we put $\mu=0$ in Theorem 1.3, we get that there exists a $g\in\mathbb{Z}[X]$, such that $|\widehat{f}(\chi)|\geq|g(\omega_{p})/p^{k}|$, where $k=\lceil\log_{p}s_{f}\rceil+1$, $s_{f}$ being the sparsity of the Boolean valued function $f$. We know by Theorem 2.1 that $|\prod_{i=1}^{p-1}g(\omega_{p}^{i})|\geq 1$. Therefore,
[…]
For $p=2$, $g(\omega_{p}^{i})$ is an integer, so
$\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}\;\geq\;\frac{1}{4s_{f}}$ (2.1)
For $p\neq 2$, the conjugate of $g(\omega_{p}^{i})$, namely $\overline{g(\omega_{p}^{i})}$, is nothing but $=g(\omega_{p}^{p-i})$. Since $|g|_{1}\leq p^{k}$, it follows that for any $i\in[p-1]$, $|g(\omega_{p}^{i})/p^{k}|\leq 1$. Therefore,
[…]
$\Rightarrow\prod_{i=1}^{(p-1)/2}\bigg{|}\frac{g(\omega_{p}^{i})}{p^{k}}\bigg{|}^{2}\;\geq\;(\frac{1}{p^{k}})^{p-1}$
$\Rightarrow\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}^{2}\;\geq\;\frac{1}{p^{k(p-1)}}$
$\Rightarrow\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}\;\geq\;\frac{1}{p^{k(p-1)/2}}\geq\frac{1}{(p^{2}s_{f})^{(p-1)/2}}\;.$ (2.2)
So, from Equation 2.1 and 2.2, we have
$\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}\;\geq\;\frac{1}{(p^{2}s_{f})^{\lceil(p-1)/2\rceil}}\;.$
Remark 2.3.
Let $p$ be a prime. The proof-technique of Theorem 1.5 gives the following. If the Fourier coefficients of a Boolean function $f$ are $k$ granular, where $k=\lceil\log_{p}s\rceil+1$, then the Fourier coefficients of $f^{2}$ are $2k$-granular, and their absolute values are $\geq\frac{1}{(p^{2}s)^{(p-1)}}$.
3 Sparse Boolean-valued function with small Fourier coefficients
In this section, for a fixed prime $p\geq 5$, and arbitrarily large $s$, we give an example of a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, such that the minimum of the absolute value of its Fourier coefficients is at most $o\left(1/s\right)$. In this section, we give the details of the construction of Theorem 1.6; the details are presented in the Section D.1 and Section D.2.
To prove Theorem 1.6 we define a function, which is basically composition of $\operatorname{AND}_{n}$ and univariate Threshold functions; we call $\operatorname{\mathsf{AT}}$.
Definition 3.1.
Let us define function $\operatorname{\mathsf{AT}}:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ by
$\operatorname{\mathsf{AT}}(x_{1},x_{2},\ldots,x_{n})\;:=\;\operatorname{AND}_{n}\left(\mathbb{I}_{\geq\frac{p+1}{2}}(x_{1}),\mathbb{I}_{\geq\frac{p+1}{2}}(x_{2}),\ldots,\mathbb{I}_{\geq\frac{p+1}{2}}(x_{n})\right),$
where the univariate Threshold function $\mathbb{I}_{\geq\frac{p+1}{2}}:\mathbb{Z}_{p}\to\{-1,+1\}$, is defined as:
$\mathbb{I}_{\geq\frac{p+1}{2}}(x)\;=\;\begin{cases}1&\text{ for }x\geq\frac{p+1}{2}\\ -1&\text{ otherwise.}\end{cases}$
Lemma 3.2.
There is a Fourier coefficient of $\operatorname{\mathsf{AT}}$, whose absolute value is $\frac{1}{p^{nc}}$, where $c$ is a constant $>1$.
The proof of Lemma 3.2 is given in Section D.1 due to lack of space.
Now we are ready to prove Theorem 1.6.
Proof 3.3 (Proof of Theorem 1.6).
This directly follows, as we can claim from Lemma 3.2 that there exists a family of functions in $\mathbb{Z}_{p}^{n}$ whose absolute value of the minimum coefficient is not linear in $\frac{1}{\text{sparsity}}$, but actually $=\Omega(\frac{1}{\text{sparsity}^{1+\epsilon_{p}}})$. where $\epsilon_{p}>0$, is a $p$-dependent constant.
Lemma 3.4.
There exists a function whose one of the Fourier coefficients is less than that of $\operatorname{\mathsf{AT}}$.
The proof of Lemma 3.4 is given in Section D.2 due to lack of space.
4 Testing Algorithm for Sparsity
Here we present the whole algorithm (Algorithm 1) for sparsity testing over $\mathbb{Z}_{p}^{n}$ (see next page). However, we describe the various steps of the algorithm in Section E.1, including the correctness and the analysis of the algorithm in Section E.1 except for the proofs of two crucial lemmas that are proved in Section E.2 and Section E.3.
Algorithm 1 Test Sparsity
1:  Input: $s,\epsilon$, and query access to $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$.
2:  Output: YES, if $f$ is $s$-sparse, and NO, if it is $\epsilon$-far from any Boolean valued function.
3:  Parameter setting: Set $t:=\lceil 2\log_{p}s+\log_{p}20\rceil+1,\tau:=\min(\frac{\epsilon^{2}}{40p^{t}},\frac{1}{(p^{2}s)^{p-1}})$ $M:=O(\log(p^{t})\cdot\frac{1}{\tau^{2}})$
4:  Choose $v_{1},\dots,v_{t}$ linearly independent vectors uniformly at random from $\mathbb{Z}_{p}^{n}$.
5:  Let $H=\operatorname{Span}\{v_{1},\dots,v_{t}\}$
6:  Pick $(z_{1},x_{1}),\dots,(z_{M},x_{M})$ uniformly and independently from $H\times\mathbb{Z}_{p}^{n}$
7:  Query $f(x_{1}),\dots,f(x_{M})$ and $f(x_{1}-z_{1}),\dots,f(x_{M}-z_{M})$
8:  for For every $r\in\mathbb{Z}_{p}^{n}$ do
9:     Let $\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$ be the estimate of $wt(r+H^{\perp})$
10:  end for
11:  if number of $r$ for which the estimate of $wt(r+H^{\perp})$ is $\geq\frac{2\tau}{3}$ is $\leq s$ then
12:     Output YES
13:  else
14:     Output NO
15:  end if
5 Conclusion
Gopalan et al. [15] was the first to study the problem of testing Fourier sparsity of Boolean function over $\mathbb{Z}^{n}_{2}$. Along the way, they were able to drive fundamental properties of Boolean functions over $\mathbb{Z}_{2}^{n}$, like Granularity of the Fourier spectrum, that have found many other applications [4, 8]. In this work, we have extended their results for groups that can be written as $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$.
The most natural question that arises from our work will be to study this problem for (finite) general Abelian groups. Unfortunately, our work does not extend to finite Abelian groups, because of the absence of the component-wise vector space structure. In particular, it would be nice to get a lower bound on the absolute value of non-zero Fourier coefficients of a sparse Boolean-valued function over $\mathbb{Z}_{p^{2}}^{n}$. In any way, our results in this paper should be seen as a first step toward solving the lower bound problem over finite Abelian group.
Finally, we ask whether it is possible to show a better ($p$-dependent) lower bound over $\mathbb{Z}_{p}^{n}$ on the query-complexity of any adaptive sparsity testing algorithm.
Appendix A Preliminaries
A.1 Fourier Analysis over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$
$\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ forms a finite Abelian group under addition whose order is $p_{1}^{n_{1}}\cdots p_{t}^{n_{t}}$, where $p_{1},\ldots,p_{t}$ are distinct primes. Its individual components form a vector space, that is, $\mathbb{Z}_{p_{i}}^{n_{i}}$ is a vector space over the field $\mathbb{Z}_{p_{i}}$ for all $i\in\{1,2,\ldots t\}$ whenever $p_{i}$’s are primes. Throughout this paper we will denote $\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ by $\mathcal{G}$. So $\mathcal{G}$ is a finite Abelian group with $|\mathcal{G}|=p_{1}^{n_{1}}\cdots p_{t}^{n_{t}}$, where |.||.|| . | denotes the order of $\mathcal{G}$.
We will denote this root of unity by $\omega_{p}$ a $p^{th}$ primitive root of unity, that is $e^{2\pi\iota/p}$.
Let us begine by defining the characters of $\mathcal{G}$.
Definition A.1.
(Character) A character of $\mathcal{G}$ is a homomorphism $\chi:\mathcal{G}\to\mathbb{C}^{\times}$ of $\mathcal{G}$, that is, $\chi$ satisfies the following:
$\chi(x+y)=\chi(x)\chi(y),\ x,y\in\mathcal{G}.$
Equivalently, a character $\chi$ of $\mathcal{G}$ can be defined by
$\chi(x_{1},\ldots,x_{t})=\chi_{r_{1}}(x_{1})\cdots\chi_{r_{t}}(x_{t})=\omega_{p_{1}}^{r_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{r_{t}\cdot x_{t}},$
where $\chi_{r_{i}}$ is a character of $\mathbb{Z}_{p_{i}}^{n_{i}}$ for each $i$ and is defined by $\chi_{r_{i}}(x_{i})=\omega_{p_{i}}^{r_{i}\cdot x_{i}}$, $x_{i},r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$, $r_{i}\cdot x_{i}=\sum_{j=1}^{n_{i}}r_{i(j)}x_{i(j)}$, $r_{i(j)}$ and $x_{i(j)}$ being the $j^{th}$ component of $r_{i}$ and $x_{i}$ respectively (It follows from the fact that any character $\chi$ of $\mathcal{G}$ can be written as the product of characters of $\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,\mathbb{Z}_{p_{t}}^{n_{t}}$). Let us denote this character by $\chi_{r_{1},\ldots,r_{t}}$.
Now let us look at some properties of characters.
Lemma A.2.
Let $\chi$ be a character of $\mathcal{G}$. Then,
  1. 1.
    $\chi_{0}(x)=1$ for all $x\in\mathcal{G}$.
  2. 2.
    $\chi(-x)=\chi(x)^{-1}=\overline{\chi(x)}$ for all $x\in\mathcal{G}$.
  3. 3.
    For any character $\chi$ of $\mathcal{G}$, where $\chi\neq\chi_{0}$, $\sum_{x\in\mathcal{G}}\chi(x)=0$.
  4. 4.
    $|\chi_{0}(x)|=1$ for all $x\in\mathcal{G}$.
Now let us define the dual group of $\mathcal{G}$.
Definition A.3.
(Dual group) The set of characters of $\mathcal{G}$ forms a group under the operation $(\chi\psi)(x)=\chi(x)\psi(x)$ and is denoted by $\widehat{\mathcal{G}}$, where $\chi$ and $\psi$ are characters of $\mathcal{G}$. $\widehat{\mathcal{G}}$ is called the dual group of $\mathcal{G}$.
The following theorem states that $\mathcal{G}$ is isomorphic to its dual group.
Theorem A.4.
$\widehat{\mathcal{G}}\cong\mathcal{G}$.
Let us look at the definition of Fourier transform for functions on $\mathcal{G}$.
Definition A.5.
(Fourier transform) For any function $f:\mathcal{G}\to\mathbb{C}$, the Fourier transform $\widehat{f}:\widehat{\mathcal{G}}\to\mathbb{C}$ is defined by
$\widehat{f}(\chi_{r_{1},\ldots,r_{t}})=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\omega_{p_{1}}^{-r_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{-r_{t}\cdot x_{t}},$
where $x_{i},r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$, and $r_{i}\cdot x_{i}=\sum_{j=1}^{n_{i}}r_{i(j)}x_{i(j)}$, $r_{i(j)}$ and $x_{i(j)}$ being the $j^{th}$ component of $r_{i}$ and $x_{i}$ respectively.
Remark A.6.
The Fourier transform of a function $f:\mathcal{G}\to\mathbb{C}$ is defined by
$\widehat{f}(\chi)=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)},$
where $\overline{\chi(x)}$ is the conjugate of $\chi(x)$. The Definition A.5 follows from this, as $\chi=\chi_{r_{1},\ldots,r_{t}}$ for some $r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}},\ i\in\{1,\ldots,t\}$.
The following theorem states that any function from $\mathcal{G}$ to $\mathbb{C}$ can be written as a linear combination of characters of $\mathcal{G}$.
Theorem A.7.
(Fourier inversion formula) Any function $f:\mathcal{G}\to\mathbb{C}$ can be uniquely written as a linear combination of characters of $\mathcal{G}$, that is,
$f(x)=\sum_{\chi_{r_{1},\ldots,r_{t}}\in\widehat{\mathcal{G}}}\widehat{f}(\chi_{r_{1},\ldots,r_{t}})\omega_{p_{1}}^{r_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{r_{t}\cdot x_{t}},$
where $x_{i},r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$, $r_{i}\cdot x_{i}=\sum_{j=1}^{n_{i}}r_{i(j)}x_{i(j)}$, $r_{i(j)}$ and $x_{i(j)}$ being the $j^{th}$ component of $r_{i}$ and $x_{i}$ respectively.
Theorem A.8.
(Parseval) For any two functions $f,g:\mathcal{G}\to\mathbb{C}$,
$\mathbb{E}_{x\in\mathcal{G}}[f(x)\overline{g(x)}]=\sum_{\chi\in\widehat{\mathcal{G}}}\widehat{f}(\chi)\overline{\widehat{g}(\chi)}.$
More specifically, if $f:\mathcal{G}\to\{-1,+1\}$ is a Boolean-valued function then
$\sum_{\chi\in\widehat{\mathcal{G}}}|\widehat{f}(\chi)|^{2}=1.$
Now let us define the Fourier sparsity of a function $f$ on $\mathcal{G}$.
Definition A.9.
(Sparsity and Fourier Support)
  • The Fourier sparsity $s_{f}$ of a function $f:\mathcal{G}\to\mathbb{C}$ is defined to be the number of non-zero Fourier coefficients in the Fourier expansion of $f$ (Theorem A.7). In this paper, by sparsity of a function, we will mean the Fourier sparsity of the function.
  • Fourier support $\operatorname{supp}(f)$ of a function $f:\mathcal{G}\to\mathbb{C}$ denotes the set $\left\{\chi\,\mid\,\widehat{f}(\chi)\neq 0\right\}$.
The following lemma states an important property of characters of a subgroup $H=H_{1}\times\cdots\times H_{t}$ of $\mathcal{G}$.
Lemma A.10.
$\sum_{h\in H}\chi_{h}=|H|\cdot 1_{H^{\perp}}$, where $H=H_{1}\times\cdots\times H_{t}$ is a subgroup of $\mathcal{G}$, and $h=(h_{1},\ldots,h_{t}),\ h_{i}\in H_{i}\ \forall i$. That is,
$\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\chi_{h_{1}}\cdots\chi_{h_{t}}=|H_{1}|\cdots|H_{t}|\cdot 1_{H_{1}^{\perp}}\cdots 1_{H_{t}^{\perp}}.$
Here, for each $i$, $H_{i}$ is a subgroup of $\mathbb{Z}_{p_{i}}^{n_{i}}$ and hence a subspace $\mathbb{Z}_{p_{i}}^{n_{i}}$, as $\mathbb{Z}_{p_{i}}^{n_{i}}$ is a vector space.
Proof A.11.
For each $i\in\{1,\ldots,t\}$, we know that for $h_{i}\in H_{i}$, $h_{i}\cdot x=0\ \forall x\in H_{i}^{\perp}$. Also, since $H_{i}$ is a subspace of $\mathbb{Z}_{p_{i}}^{n_{i}}$ for each $i$ (hence a subgroup of $\mathbb{Z}_{p_{i}}^{n_{i}}$), we have the characters of $H_{i}$ to be the induced characters of $\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$. Hence,
$\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\chi_{h_{1}}\cdots\chi_{h_{t}}(x_{1},\ldots,x_{t})=0$
for all $(x_{1},\ldots,x_{t})\notin(H_{1}^{\perp},\ldots,H_{t}^{\perp})$ by part (3) of Lemma A.2. Now, for $(x_{1},\ldots,x_{t})\in(H_{1}^{\perp},\ldots,H_{t}^{\perp})$ we have,
$\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\chi_{h_{1}}(x_{1})\cdots\chi_{h_{t}}(x_{t})$ $=\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\omega_{p_{1}}^{h_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{h_{t}\cdot x_{t}}$
$=\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}1,\text{ since }h_{i}\cdot x_{i}=0\text{ for all }i$
$=|H_{1}|\cdots|H_{t}|.$
Hence, $\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\chi_{h_{1}}\cdots\chi_{h_{t}}=|H_{1}|\cdots|H_{t}|\cdot 1_{H_{1}^{\perp}}\cdots 1_{H_{t}^{\perp}}$.
Lemma A.12.
Let $f,g$ be two Boolean valued functions from $\mathcal{G}$ to $\{-1,+1\}$. Then,
$|\widehat{fg}(\chi)|\leq||f||_{2}||g||_{2},$
for any character $\chi\in\widehat{\mathcal{G}}$.
Proof A.13.
$|\widehat{fg}(\chi)|$ $=|\sum_{\psi\in\widehat{\mathbb{Z}_{p}^{n}}}\widehat{f}(\psi)\widehat{g}(\chi+\psi)|$
[…]
Now let us formally define the notion of $\epsilon$-close and $\epsilon$-far in $\ell_{2}$ below.
Definition A.14 ($\mu$-close to $s$-sparse).
Let $f$ and $g$ be two functions with domain $\mathbb{Z}_{p}^{n}$ and range $\mathbb{C}$. Then the square of the $\ell_{2}$ distance between $f$ and $g$ is defined as $\mathbb{E}_{x\in\mathcal{G}}[|f(x)-g(x)|^{2}].$ By Parseval’s identity the square of the $\ell_{2}$-distance between $f$ and $g$ can also be written as $\sum_{\chi\in\widehat{\mathbb{Z}_{p}^{n}}}|\widehat{(f-g)}(\chi)|^{2}$.
We say that $f$ is $\epsilon$-close to $g$ in $\ell_{2}$ if the square of the $\ell_{2}$ distance between $f$ and $g$ is less than $\epsilon$. Similarly, $f$ is $\epsilon$-far from $g$ in $\ell_{2}$ if the square of the $\ell_{2}$ distance between $f$ and $g$ is at least $\epsilon$.
The following lemma gives us a relation between the $\ell_{2}$ distance between two Boolean valued functions $f$ and $g$ defined in Definition A.14 and $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.
Lemma A.15.
The square of the $\ell_{2}$ distance between two Boolean valued functions $f$ and $g$ defined in Definition A.14 is equal to $4\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.  [^2] If $f$ and $g$ are two Booelan-valued functions then $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$ is also called the Hamming distance between the two functions. So the $\ell_{2}$ norm between two Boolean-valued functions is 4 times the Hamming distance between two Boolean-valued functions.
Proof A.16.
By Parseval (Theorem A.8), we have,
$\mathbb{E}_{x\in\mathcal{G}}[(f-g)(x)\overline{(f-g)(x)}]$ $=\sum_{\chi\in\widehat{\mathcal{G}}}|\widehat{(f-g)}(\chi)|^{2}.$
Also,
$\mathbb{E}_{x\in\mathcal{G}}[(f-g)(x)\overline{(f-g)(x)}]$ $=\frac{1}{|\mathcal{G}|}\times(4\times|\{x\in\mathcal{G}|(f-g)(x)\neq 0\}|)$
$=\frac{1}{|\mathcal{G}|}\times(4\times|\{x\in\mathcal{G}|f(x)\neq g(x)\}|)$
$=4\times\Pr_{x}[f(x)\neq g(x)],$
where $|S|$ denotes the number of elements in the set $S$.
So, the square of the $\ell_{2}$ distance between $f$ and $g$ is equal to $4$ times $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.
Now let us define the total variation distance between two probability distributions.
Definition A.17.
Let $(\Omega,\mathcal{F})$ be a probability space, and $P$ and $Q$ be probability distributions defined on $(\Omega,\mathcal{F})$. The total variation distance between $P$ and $Q$ is defined in the following way.
$||P-Q||_{TV}=\sup_{A\in\mathcal{F}}|P(A)-Q(A)|.$
Lemma A.18.
Given two probability distributions $P$ and $Q$ on a probability space $(\Omega,\mathcal{F})$, the total variation distance between $P$ and $Q$ is half of the $L_{1}$ distance between them. That is,
$||P-Q||_{TV}=\frac{1}{2}\sum_{x}|P(x)-Q(x)|.$
A.2 Basics from Probability
We will need some notions and lemmas from probability. We will need the following notion of covariance between two random variables.
Definition A.19 (Covariance).
Let $X_{1}$ and $X_{2}$ be two random variables. Then, the covariance of $X_{1}$ and $X_{2}$ is given by
$\mathsf{cov}(X_{1},X_{2})=\mathbb{E}[X_{1}X_{2}]-\mathbb{E}[X_{1}]\mathbb{E}[X_{2}].$
We will need the following concentration lemmas. Let us look at Markov’s inequality first.
Lemma A.20 (Markov’s inequality).
Let $X$ be a nonnegative random variable and $a>0$, then,
$\Pr[X\geq a]\leq\frac{\mathbb{E}[X]}{a},$
where $\mathbb{E}[X]$ is the expected value of $X$.
Let us now state Chernoff’s bound for random variables taking value in [-1, 1].
Lemma A.21.
(Chernoff bound) Let $X_{1},\ldots,X_{n}$ be independent random variables taking values in $\{0,1\}$. Let $X=\sum_{i=1}^{n}X_{i}$ and $\mathbb{E}[X]$ be the expected value of $X$. Then,
  1. 1.
    $\Pr[X\leq(1-\delta)\mathbb{E}[X]]\leq e^{-\frac{\delta^{2}\mathbb{E}[X]}{2}}$, $10<\delta<$ 1.
  2. 2.
    $\Pr[X\geq(1+\delta)\mathbb{E}[X]]\leq e^{-\frac{\delta^{2}\mathbb{E}[X]}{2+\delta}}$, $\delta\geq 0$.
Now, we state the Hoeffding’s inequality for random variables taking value in [-1, 1].
Lemma A.22 (Hoeffding’s Inequality).
Let $X_{1},\cdots,X_{k}$ be real independent random variables, each taking value in [-1, 1]. Then
$\Pr\left[\left|\sum_{i=1}^{k}X_{i}-\mathbb{E}\sum_{i=1}^{k}X_{i}\right|\geq\epsilon\right]\leq 2\exp(-\frac{\epsilon^{2}}{2k})$
A.3 Basics from algebraic number theory
Fix $n\in\mathbb{N}$, and let us denote $e^{2\pi\iota/n}$, a primitive $n$-th root of unity, by $\omega_{n}$. Let $\mathbb{Q}(\omega_{n})$, be the field obtained by adjoining $\omega_{n}$ to $\mathbb{Q}$, the field of rationals. One can think of $\mathbb{Q}(\omega_{n})=\{r(\omega_{n})\mid r\in\mathbb{Q}[x]\}$, where $\mathbb{Q}[x]$ is the ring of polynomials with coefficients from $\mathbb{Q}$. Similarly, one can define $\mathbb{Z}[\omega_{n}]$, by adjoining $\omega_{n}$ to the domain of integers $\mathbb{Z}$.
Definition A.23.
(Cyclotomic Integer) Any $\alpha\in\mathbb{C}$ is called a cyclotomic integer, if $\alpha\in\mathbb{Z}[\omega_{n}]$.
Recall that $\alpha\in\mathbb{C}$ is an algebraic number if it is the root of a non-zero polynomial in $\mathbb{Q}[x]$. The minimal polynomial of $\alpha$ is the unique monic polynomial in $\mathbb{Q}[x]$ having $\alpha$ as a root. If the minimal polynomial has integer coefficients, then we say that $\alpha$ is an algebraic integer. It is a well-known fact that a cyclotomic integer is an algebraic integer. For details, see [23, Chapter 13], or expository lecture notes [16, Chapter 2] & [30].
Recall that the Galois group, denoted $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$ of automorphisms of $Q(\omega_{n})$; it is isomorphic to the multiplicative group $\mathbb{Z}_{n}^{\ast}$, of integers $\bmod\leavevmode\nobreak\ n$, i.e., all integers $k$ such that gcd$(k,n)=1$. Note that $|\mathbb{Z}_{n}^{\ast}|=\varphi(n)$, where $\varphi$ is the Euler’s totient function, where $\varphi(n):=n\cdot\prod_{\text{prime}\leavevmode\nobreak\ p\mid n}(1-1/p)$. In simple words, think of $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$, as consisting $\varphi(n)$ many automorphism $\sigma_{k}$, that sends $\omega_{n}\mapsto\omega_{n}^{k}$, where $k\in\mathbb{Z}_{n}^{\ast}$. For basics, we refer to [16, Chapter 1].
Finally, as follows, one can define norm of an $\alpha\in\mathbb{Q}(\omega_{n})$.
Definition A.24 (Norm).
The norm of $\alpha\in\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$ is defined by
$N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q}}(\alpha)\;:=\;\prod_{\sigma\in\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})}\,\sigma(\alpha)\;.$
It is a well-known fact that for a non-zero $\alpha$, $N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q}}(\alpha)\in\mathbb{Q}-\{0\}$; see [16, Lemma 1.6.1]. In particular, when $\alpha$ is a cyclotomic integer, i.e, $\alpha\in\mathbb{Z}[\omega_{n}]$, it is well-known that $N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q}}(\alpha)\in\mathbb{Z}$, see [16, Corollary 2.2.2]. We restate the same as a theorem below in the most simplified way without the above terminologies since this fact has been crucially used throughout the paper.
Theorem A.25.
For $n\in\mathbb{Z}$, let $\omega_{n}$ be a primitive root of unity. Let $f\in\mathbb{Z}[x]$, such that $f(\omega_{n})\neq 0$. Then,
[…]
In retrospect, one can think of $\alpha=f(\omega_{n})\in\mathbb{Z}[\omega_{n}]$. Note that, $\sigma_{k}(\alpha)=f(\omega_{n}^{k})$, by the definition of the automorphism $\sigma:\omega_{n}\mapsto\omega_{n}^{k}$, where gcd$(k,n)=1$. Therefore, their product, which is the norm $N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})}(\alpha)$, must be a non-zero integer, and hence the conclusion follows. Using Theorem 2.1, one can directly conclude the following fact (which can also be proved in an elementary way).
Fact 1.
Let $f\in\mathbb{Z}[x]$, such that $f(\omega_{p})\neq 0$. Then, $f(\omega_{p}^{i})\neq 0$, for any $i\in[p-1]$.
Appendix B Proofs of structure Theorems 1.3 and  1.4
In this section, we will prove Theorem 1.3 and Theorem 1.4 which are generalizations of Gopalan et al’s result  [15] to the case when the domain of the function is $\mathbb{Z}_{p}^{n}$, where $p$ is an odd prime. Let $\omega_{p}$ be a primitive $p^{th}$ root of unity. For any polynomial $\sum_{i=0}^{d}g_{i}x^{i}=g(x)\in\mathbb{C}[x]$, let $|g|_{1}\;:=\;\sum_{i=0}^{d}|g_{i}|$.
The proof of the more general result of Theorem 1.5 has been relegated to the Appendix F.
B.1 Proof of Theorem 1.3
In this subsection we will prove Theorem 1.3. Recall the definition of granularity for complex numbers (Definition 1.1) and the definition of $\mu$-close to granular (Definition 1.2). Before moving forward, we need the following definition.
Definition B.1.
(Projection operator $P_{r+H}$) Given a subgroup $H=H_{1}\times\cdots\times H_{t}\subseteq\mathcal{G}$ and a coset $r+H$, where for each $i$, $H_{i}$ is a subgroup of $\mathbb{Z}_{p_{i}}^{n_{i}}$ (hence a subspace), let us define the projection operator $P_{r+H}$ on functions $f:\mathcal{G}\to\mathbb{R}$ as follows:
$\widehat{P_{r+H}f}(\chi)=\begin{cases}\widehat{f}(\chi)&\chi\in r+H\\ 0&\text{otherwise}.\end{cases}$
So,
$P_{r+H}f(x)=\sum_{\chi\in r+H}\widehat{f}(\chi)\chi(x).$
To prove Theorem 1.5, we need to prove Theorem 1.3 and to prove Theorem 1.3, we need the following lemma.
Lemma B.2.
Given a subspace $H$ of $\mathbb{Z}_{p}^{n}$, we have
$P_{r+H}f(x)\;=\;\mathbb{E}_{z\in H^{\perp}}[f(x-z)\chi_{r}(z)]\;=\;\frac{1}{|H^{\perp}|}\sum_{z\in H^{\perp}}f(x-z)\chi_{r}(z)\;.$
Proof B.3 (Proof of Lemma B.2).
From the definition of the projection operator $P_{r+H}$ (Definition B.1) we get
$P_{r+H}f(x)$ $=\sum_{\alpha\in r+H}\hat{f}(\alpha)\chi_{\alpha}(x)$
[…]
$=\sum_{\beta\in H}\hat{f}(r+\beta)\chi_{r}(x)\chi_{\beta}(x)$
$=\chi_{r}(x)\sum_{\beta\in H}\hat{f}(r+\beta)\chi_{\beta}(x)$
$=\chi_{r}(x)\sum_{\beta\in H}\chi_{\beta}(x)\frac{1}{|\mathbb{Z}_{p}^{n}|}\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\overline{\chi_{r+\beta}(y)}$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\sum_{\beta\in H}\chi_{\beta}(x)\chi_{-r-\beta}(y)$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\sum_{\beta\in H}\chi_{\beta}(x)\chi_{r+\beta}(-y)$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\chi_{r}(-y)\sum_{\beta\in H}\chi_{\beta}(x-y)$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\chi_{r}(-y)|H|1_{H^{\perp}}(x-y)$
$=\frac{1}{|H^{\perp}|}\chi_{r}(x)\sum_{x-y\in H^{\perp}}f(y)\chi_{r}(-y)$
$=\frac{1}{|H^{\perp}|}\sum_{x-y\in H^{\perp}}f(y)\chi_{r}(x-y)$
[…]
$=\mathbb{E}_{z\in H^{\perp}}[f(x-z)\chi_{r}(z)].$
Now we prove Theorem 1.3.
Proof B.4 (Proof of Theorem 1.3).
Set $k=\lceil\log_{p}s\rceil+1$, and let $A$ be a $k\times n$ matrix and $b$ be a $k$-dimensional vector.
Let $A^{\perp}\subset\mathbb{Z}^{n}_{p}$, be the set of solutions to $A\chi=0$. Let $H$ be the coset of $A^{\perp}$ that are solutions to system of linear equations $A\chi=b$. From the definition,
$P_{H}f(x)\;=\;\sum_{\chi\in H}\widehat{f}(\chi)\chi(x).$ (B.1)
We need to show that for each $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix $A$ and a $k$-dimensional vector $b$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A\chi\,=\,b\right\}\;=\;\{\chi_{r_{i}}\},$ (B.2)
and, the $\ell_{2}$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, that is,
$\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\;\leq\;\frac{\mu^{2}}{s}\,,$
where $B$ is the set of character functions corresponding to the $s$ largest coefficients and $S$ is the rest of the characters, and $H$ is the solution set of the system of linear equations $A\chi=b$. Let $E_{1}$ and $E_{2}$ be the first and the second events respectively. So we need to prove that both these events occur simultaneously with non-zero probability.
To show the existence of $A$ and $b$, given $\chi_{r_{i}}\in B$, for which Equation B.2 is true we will need to consider two cases:
Case 1: $\chi_{r_{i}}=0$
 [^3] Here $\chi_{r_{i}}=0$ means that $\chi_{r_{i}}=\chi_{r}$ where $r=(0,\dots,0)\in\mathbb{Z}^{n}_{p}$.
. Observe that for all $A$, we have $A\chi_{r_{i}}=0$. If entries of $A$ are sampled independently and uniformly at random from $\mathbb{Z}_{p}$. Then, for all $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$, we have
$\Pr_{A}\left[A\chi_{r_{j}}=0\right]=\frac{1}{p^{k}}\leq\frac{1}{s}.$
Note that $\Pr_{A}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when $A$ is the random matrix generated by the above-mentioned procedure. Using union bound, we get
$\Pr_{A}\left[\exists j\neq i\text{ such that }A\chi_{r_{j}}=0\right]\leq\frac{s-1}{s}\leq\frac{p^{k-1}-1}{p^{k}}<1$
Therefore, if $\chi_{r_{i}}=0$ then there exists a $k\times n$ matrix $A$ such that
[…]
Case 2: $\chi_{r_{i}}\neq 0$
 [^4] Here $\chi_{r_{i}}=0$ means that $\chi_{r_{i}}=\chi_{r}$ where $r\neq(0,\dots,0)\in\mathbb{Z}^{n}_{p}$.
. Fix a nonzero vector $b\in\mathbb{Z}^{n}_{p}\setminus\{(0,\dots,0)\}$. Again, let $A$ be a random $k\times n$ matrix with entries of $A$ being sampled independently and uniformly at random from $\mathbb{Z}_{p}$. Observe that $\Pr_{A}[A\chi_{r_{i}}=b]=\frac{1}{p^{k}}$. Let $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$. If $\chi_{r_{j}}=\alpha\chi_{r_{i}}$, where $\alpha\in\mathbb{Z}_{p}\setminus\{1\}$, then
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b]=0.$
Now we will consider the case when $\chi_{r_{j}}$ is not a scalar multiple of $\chi_{r_{i}}$. Observe that
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b]$ $=\frac{\Pr_{A,b}[A\chi_{r_{j}}\,=\,b,A\chi_{r_{i}}\,=\,b]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}$
$=\frac{\Pr_{A}\left[\bigcap\limits_{\ell\in[k]}\left(A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right)\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}$
$=\frac{\prod\limits_{\ell\in[k]}\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}$ (B.3)
Note that $A_{\ell}$ and $b_{\ell}$ denote the $\ell$-th row of $A$ and $\ell$-th coordinate of $b$ respectively, and $[k]$ denotes the set $\{1,\dots,k\}$. Observe that the last equality in the above equation follows from the fact that rows of $A$ are independent and uniformly random samples of $\mathbb{Z}^{n}_{p}$. As $\chi_{r_{j}}$ and $\chi_{r_{i}}$ are not scalar multiple of each other therefore
$\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right]=\frac{1}{p^{2}}.$
Using the above equation and the fact that $\Pr_{A}[A\chi_{r_{i}}=b]=\frac{1}{p^{k}}$ in Equation (Case 2: $\chi_{r_{i}}\neq 0$) we get
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b]=\frac{\prod\limits_{\ell\in[k]}\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}=\frac{\left(\frac{1}{p^{2}}\right)^{k}}{\frac{1}{p^{k}}}=\frac{1}{p^{k}}$
Using the conditional variant of the union bound we get
$\Pr_{A}[\exists j\neq i\text{ such that }A\chi_{r_{j}}=b\mid A\chi_{r_{i}}=b]\leq\frac{s-1}{p^{k}}\leq\frac{p^{k-1}-1}{p^{k}}<1.$
From the above two cases we can now say that given any $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix $A$ and a $k$-dimensional vector $b$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A\chi\,=\,b\right\}=\{\chi_{r_{i}}\}.$
Now, let us look at the coefficients corresponding to the characters in $S$, in order to prove that the $\ell$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, $H$ being the solution set of the system of linear equations $A\chi=b$. Let us look at the expected value of $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$, given $A\chi_{r_{i}}=b$.
$\mathbb{E}_{A,b}\biggl{[}\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\mid A\chi_{r_{i}}=b\biggr{]}=\sum_{\chi\in S}\Pr_{A,b}[\chi\in H\mid A\chi_{r_{i}}=b]|\widehat{f}(\chi)|^{2}\leq\frac{\mu^{2}}{p^{k}}\leq\frac{\mu^{2}}{ps},$ (B.4)
since $k=\lceil\log_{p}s\rceil+1$.
Therefore, by Markov’s inequality, we have,
$\Pr_{A,b}\bigg{[}\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\geq\frac{\mu^{2}}{s}\mid A\chi_{r_{i}}=b\bigg{]}\leq\frac{\frac{\mu^{2}}{ps}}{\frac{\mu^{2}}{s}}=\frac{1}{p}<1.$ (B.5)
So, if we apply union bound to Equation B.2 and Equation B.5, we get that the events $E_{1}$ and $E_{2}$ occur simultaneously with non-zero probability.
Therefore, we have,
$P_{H}f(x)=\widehat{f}(\chi_{r_{i}})\chi_{r_{i}}(x)+\sum_{\chi\in S\cap H}\widehat{f}(\chi)\chi(x),$
where $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\leq\frac{\mu^{2}}{s}$.
Now, by Lemma B.2, we have,
$P_{H}f(x)=\mathbb{E}_{z\in r(A)}[f(x-z)\chi_{r}(z)],$
which is of the form $\frac{g(\omega_{p})}{p^{k}}$, where $g(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq p-1$, and $r(A)$ is the row-space of $A$. So $P_{H}f(x)$ is $k$-granular.
Also, let us look at the function $G_{1}(x)=\sum_{\chi\in S\cap H}\widehat{f}(\chi)\chi(x)$. We know, by Equation B.4, that, $\mathbb{E}_{x}G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$. So, $G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$ for some $x\in\mathbb{Z}_{p}^{n}$. For this particular $x$,
$P_{H}f(x)=\widehat{f}(\chi_{r_{i}})\chi_{r_{i}}(x)+G_{1}(x).$
Therefore,
$|\widehat{f}(\chi_{r_{i}})|=|P_{H}f(x)-G_{1}(x)|,$
which implies that $\widehat{f}(\chi_{r_{i}})$ is $\frac{\mu}{\sqrt{s}}$ close to $k$-granular, since $P_{H}f(x)$ is $k$-granular and $|G_{1}(x)|\leq\frac{\mu}{\sqrt{s}}$.
B.2 Proof of Theorem 1.4
We have proved in Section B.1 that if the Fourier weight of the function $f$ is concentrated on the $s$-largest Fourier coefficients, then $f$ is $\frac{\mu}{\sqrt{s}}$ close to a $\lceil\log_{p}s\rceil+1$-granular in the $\ell_{2}$ norm. In this subsection, we will prove Theorem 1.4, which says that this function must be a Boolean valued function.
Proof B.5.
Let us assume $k=\lceil\log_{p}s\rceil+1$. Let $B$ be the set of character functions corresponding to the $s$ largest coefficients and $S$ be set of the rest of the characters. Then, by Theorem 1.3, for all $\chi_{r_{i}}\in B$, $\widehat{f}(\chi_{r_{i}})$ is $\frac{\mu}{\sqrt{s}}-$close to $\lceil\log_{p}s\rceil+1$-granular. Therefore, for all $\chi_{r_{i}}\in B$, we can write $f(\chi_{r_{i}})$ as the sum of $\widehat{F}(\chi_{r_{i}})$ and $\widehat{G}(\chi_{r_{i}})$, that is,
$\widehat{f}(\chi_{r_{i}})=\widehat{F}(\chi_{r_{i}})+\widehat{G}(\chi_{r_{i}}),$
where $|\widehat{F}(\chi_{r_{i}})|$ is $\lceil\log_{p}s\rceil+1$-granular and $|\widehat{G}(\chi_{r_{i}})|\leq\frac{\mu}{\sqrt{s}}$.
Now, for $\chi\in\mathcal{S}$, let us set $\widehat{F}(\chi)=0$ and $\widehat{G}(\chi)=\widehat{f}(\chi)$. Therefore, we can write $f(x)$ as sum of $F(x)$ and $G(x)$, that is,
$f(x)=F(x)+G(x).$
Clearly, $F$ is $s-$sparse and $|\widehat{F}(\chi)|$ is $\lceil\log_{p}s\rceil+1$-granular for all Fourier coefficients $\widehat{F}(\chi)$ of $F$.
Let us look at $G(x)$. We have,
$\mathbb{E}[G(x)^{2}]$ $\;=\;\mathbb{E}[\sum_{\chi\in\mathcal{B}}\widehat{G}(\chi)^{2}+\sum_{\chi\in\mathcal{S}}\widehat{G}(\chi)^{2}]$
$\;=\;\mathbb{E}[\sum_{\chi_{r_{i}}\in\mathcal{B}}\widehat{G}(\chi_{r_{i}})^{2}+\sum_{\chi\in\mathcal{S}}\widehat{f}(\chi)^{2}]$
$\;\leq\;s\cdot\frac{\mu^{2}}{s}+\mu^{2}$
 $\;\leq\;2\mu^{2}.$
Claim 2.
Range of $F$ is {-1,+1}.
Proof B.6 (Proof of 2).
Since $f$ is Boolean valued and $f=F+G$, so we have,
$1=f^{2}=(F+G)^{2}=F^{2}+2FG+G^{2}=F^{2}+G(2F+G)=F^{2}+G(2f-G).$ (B.6)
Let $H=G(2f-G)$. Let us look at the Fourier coefficients $\widehat{H}(\chi)$ of $H$.
$|\widehat{H}(\chi)|$ $\leq||G||_{2}||2f-G||_{2}$ by Lemma A.12
$\leq||G||_{2}(||2f||_{2}+||G||_{2})$
$\leq\sqrt{2}\mu\cdot 2\cdot 1+2\mu^{2}$
$<4\mu$
$\leq\frac{1}{2(p^{2}s)^{p-1}}.$ (B.7)
Now, since $\widehat{F}(\chi)$ is $\lceil\log_{p}s\rceil+1$-granular, any Fourier coefficient $\widehat{F^{2}}(\chi)$ of $F^{2}$ is $22\lceil\log_{p}s\rceil+$ 2-granular. That is, $\widehat{F^{2}}(\chi)=\frac{g^{\prime}(\omega_{p})}{p^{2k}}$, where $k=\lceil\log_{p}s\rceil+1$ and $g^{\prime}(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq p-1$. Also, by Remark 2.3, $|\widehat{F^{2}}(\chi)|\geq\frac{1}{(p^{2}s)^{p-1}}$.
Now, from Equation B.6, we have,
$\widehat{F^{2}}(\chi_{0})+\widehat{H}(\chi_{0})=1$
and
$\widehat{F^{2}}(\chi)+\widehat{H}(\chi)=0,\text{ for all }\chi\neq\chi_{0},$
where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathbb{Z}_{p}^{n}$, as defined in Lemma A.2, part 1.
Since $\widehat{F^{2}}(\chi)\geq\frac{1}{(p^{2}s)^{p-1}}$ and $\widehat{H}(\chi)<\frac{1}{2(p^{2}s)^{p-1}}$, these two cannot sum up to 00; therefore, $\widehat{F^{2}}(\chi)=0$, $\widehat{H}(\chi)=0$ for all $\chi\neq 0$.
Now, let us look at $\widehat{F^{2}}(\chi_{0})$. From the above, we have
[…]
$\Rightarrow$ $\,\widehat{H}(\chi_{0})\;=\;\frac{p^{2k}\,-\,g^{\prime}(\omega_{p})}{p^{2k}}\;,$
which implies that $\widehat{H}(\chi_{0})$ is $2k$-granular if $\widehat{H}(\chi_{0})\neq 0$. So, by Remark 2.3, $|\widehat{H}(\chi_{0})|\geq\frac{1}{(p^{2}s)^{p-1}}$, which is not possible since $|\widehat{H}(\chi_{0})|\leq\frac{1}{2(p^{2}s)^{p-1}}$ by B.6. So, $\widehat{H}(\chi_{0})=0$, hence $\widehat{F^{2}}(\chi_{0})=1$. Therefore $F^{2}(x)=0$ for all $x\in\mathbb{Z}_{p}^{n}$, which implies that the range of $F$ is $\{-1,+1\}$.
Now, let us find $\Pr_{x}[f(x)\neq F(x)]$. Since $F$ and $f$ both are Boolean valued, so the range of $G$ is equal to $\{-2,0,+2\}$. So,
$\Pr_{x}[f(x)\neq F(x)]$ $=\Pr_{x}[|G(x)|=2]$
$=\frac{1}{4}\mathbb{E}_{x}[G(x)^{2}]$ by Markov’s inequality
$\leq\frac{\mu^{2}}{2}.$
Therefore, the $\ell_{2}$ distance between $f$ and $F$ is $\leq\sqrt{4\times\frac{\mu^{2}}{2}}=\sqrt{2}\mu$.
Appendix C A $p$-independent lower bound
In this section, we discuss an interesting property of Boolean-valued functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$, where $p$ is a prime number, and eventually establish a $p$-independent lower bound on the Fourier coefficient of an $s$-sparse Boolean-valued function.
For $x\in\mathbb{Z}_{p}^{n}$ and $i\in[p-1]$, we define $i^{-1}x:=(i^{-1}x_{1},\ldots,i^{-1}x_{n})$, where $x=(x_{1},\ldots,x_{n})$ and $i^{-1}$ denotes the inverse of $i$ in the multiplicative group $\mathbb{Z}_{p}^{*}$ of $\mathbb{Z}_{p}$.
Lemma C.1.
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, and $\chi\in\operatorname{supp}(f)$. Then, for each $i\in\{1,2,\ldots,p-1\}$, there exists a function $h_{i}:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, defined by $h_{i}(x)=f(i^{-1}x)$ such that
  1. 1.
    $\hat{h}_{i}(\chi)=\hat{f}(\chi^{i})$,
  2. 2.
    $\hat{h}_{i}(\chi)\neq 0\iff\hat{f}(\chi)\neq 0$,
where $i^{-1}$ denotes the inverse of $i$ in the multiplicative group $\mathbb{Z}_{p}^{*}$ of $\mathbb{Z}_{p}$.
Proof C.2.
Proof of Part (1). Let $h_{i}(x):=f(i^{-1}x)$. Note that, by definition, $h_{i}$ is a Boolean-valued function. Further, by Definition A.5:
$\hat{h_{i}}(\chi)$ $=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}h_{i}(x)\overline{\chi(x)}$
$=\frac{1}{p^{n}}\sum_{ix\in\mathbb{Z}_{p}^{n}}h_{i}(ix)\overline{\chi(ix)}$
$=\frac{1}{p^{n}}\sum_{ix\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)^{i}}$
$=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}^{i}$
$=\hat{f}(\chi^{i}).$
Proof of Part (2). We want to prove the following.
Claim 3.
$\hat{h_{i}}(\chi)\neq 0\text{ iff }\hat{f}(\chi)\neq 0$.
Proof C.3.
By Definition A.5, we have $\hat{f}(\chi)=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}=\frac{u(\omega_{p})}{p^{n}}$, where $u(\omega_{p})=\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}$. Think of $u(X)\in\mathbb{Z}[X]$. Then, $\hat{h_{i}}(\chi)=\hat{f}(\chi^{i})=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}^{i}=\frac{u(\omega_{p}^{i})}{p^{n}}$; the last holds since
$u(\omega_{p})\;=\;\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}\;\iff\;u(\omega_{p}^{i})\;=\;\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}^{i}\;.$
Using Fact 1, we know that $u(\omega_{p})\neq 0\iff u(\omega_{p}^{i})\neq 0$. Hence, whenever $f(\chi)\neq 0,\ h_{i}(\chi)\neq 0$ as well, and the vice versa. Thus, the claim holds.
The following corollary follows immediately from Lemma C.1 (part 2).
Corollary C.4.
$s_{f}=s_{h_{i}}$, where $s_{h_{i}}$ is the sparsity of the function $h_{i}$.
Proof C.5.
From the above claim, one concludes that the function $h_{i}$ has the same sparsity as the function $f$, that is, $s_{h_{i}}=s_{f}$.
The following corollary follows immediately from Lemma C.1.
Corollary C.6.
For all function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, $s_{f}\geq p-1$.
Proof C.7.
From the proof of Lemma C.1, it is clear that if $\chi\in\operatorname{supp}(f)$, then $\chi^{i}\in\operatorname{supp}(f)$ for all $i\in\{1,\ldots,p-1\}$. Hence, sparsity of $f,\ s_{f}\geq p-1$.
Now, we prove a lower bound on the Fourier coefficient which is $p$-independent; this is an immediate consequence of Corollary C.6.
Theorem C.8 ($p$-independent lower bound).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a function. Let $\chi\in\sup(f)$. Then,
$|\widehat{f}(\chi)|\;>\;\frac{1}{((s_{f}+1)\sqrt{s_{f}})^{s_{f}}}\;.$
Proof C.9.
We know that $s_{f}\geq p-1$, by Corollary C.6. Hence, we combine this with Theorem 1.5 , to get
$|\widehat{f}(\chi)|\;\geq\;\frac{1}{(p^{2}s_{f})^{(p-1)/2}}\;>\;\frac{1}{((s_{f}+1)\sqrt{s_{f}})^{s_{f}}};.$
In the above, we used the fact that $(p^{2}s_{f})^{1/2}=p\sqrt{s_{f}}\leq(s_{f}+1)\sqrt{s_{f}}$.
Appendix D Proofs of Lemma 3.2 and Lemma 3.4
In this section, we prove Lemma 3.2 and Lemma 3.4, which are stated in Section 3.
D.1 Proof of Lemma 3.2
Proof D.1 (Proof of Lemma 3.2).
Let us look at the coefficient $\widehat{I_{\geq\frac{p+1}{2}}}(\chi_{2})$, where $\chi_{2}(x)=\omega_{p}^{2\cdot x}$. By definition,
$\widehat{I_{\geq\frac{p+1}{2}}}(\chi_{2})\;$ $=\;\frac{1}{p}(-1+\omega_{p}-\omega_{p}^{2}+\omega_{p}^{3}-\cdots)$
$=\;-\frac{1}{p}\cdot\frac{1+\omega_{p}^{p}}{1+\omega_{p}}=-\frac{2}{p}\cdot\frac{1}{1+\omega_{p}}\;.$
where we simply used the GP series with the ratio $-\omega_{p}$. Note that
$|1+\omega_{p}|^{2}=(1+\cos(2\pi/p))^{2}+\sin^{2}(2\pi/p)=2(1+\cos(2\pi/p))=4\cos^{2}(\pi/p)\;.$
Therefore,
$\left|\widehat{I_{\geq\frac{p+1}{2}}}(\chi_{2})\right|=\frac{1}{p\cos(\pi/p)}\;.$
We are not using mod, because for $p\geq 5$, $\cos(\pi/p)>0$. Now, let us consider the function $\operatorname{AND}_{n}:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$ defined by
$\operatorname{AND}_{n}(x_{1},x_{2},\ldots,x_{n})=\begin{cases}-1&\text{ if }x_{i}=1\text{ for all }i\\ 1&\text{otherwise}.\end{cases}$
It is already known that the coefficient of $(-1)^{x_{1}+\cdots+x_{n}}$ is $\frac{(-1)^{n-1}}{2^{n-1}}$. Now, let us look at the function
$\operatorname{\mathsf{AT}}(x_{1},x_{2},\ldots,x_{n})=\operatorname{AND}_{n}\left(\mathbb{I}_{\geq\frac{p+1}{2}}(x_{1}),\mathbb{I}_{\geq\frac{p+1}{2}}(x_{2}),\ldots,\mathbb{I}_{\geq\frac{p+1}{2}}(x_{n})\right).$
We know that other than the constant coefficient, the absolute value of all other coefficients of $\operatorname{AND}_{n}$ is $\frac{1}{2^{n-1}}$. So, in particular,
[…]
Comparing $p^{nc}$ with the above, one gets that
[…]
Note that, for large enough $n$, and a fixed $p\geq 5$, $c>1$. Hence, the proof follows.
D.2 Proof of Lemma 3.4
Proof D.2 (Proof of Lemma 3.4).
Consider the function $f:\mathbb{Z}_{5}^{2}\to\{-1,+1\}$ whose truth table is given in Table 1.
Observe that:
  • $\widehat{f}(\chi_{1,0})=\frac{1}{25}(-5+5\omega_{5}-5\omega_{5}^{2}+\omega_{5}^{3}+\omega_{5}^{4})=\frac{0.29}{25}.$
  • $s_{f}=25$, where $s_{f}$ is the sparsity of $f$.
Therefore,
$\widehat{f}(\chi_{1,0})=\frac{1}{s_{f}^{1+0.385}}.$
So comparing with Equation D.1, we can see that $c$ is bigger in this case.
Remark D.3.
When $p=3$, we know that $12\cos(\pi/3)=$ 1, and hence $c<1$ ($c\to 1$, when $n\to\infty$).
Remark D.4.
When $p=5$, and $n=2$, we have,
$c=\log_{5}|(2\cos(\frac{\pi}{5}))|+1-\frac{1}{2}\cdot\log_{5}(2)=1.265...$ (D.1)
Table 1: Truthtable for $f:\mathbb{Z}^{2}_{5}\rightarrow\{-1,+1\}$.
$x_{1}$ $x_{2}$ $f(x_{1},x_{2})$
0 0    -1
0 1     1
0 2    -1
0 3     1
0 4    -1
1 0    -1
1 1     1
1 2    -1
1 3     1
1 4    -1
2 0    -1
2 1     1
2 2    -1
2 3     1
2 4    -1
3 0    -1
3 1     1
3 2    -1
3 3    -1
3 4    -1
4 0    -1
4 1     1
4 2    -1
4 3    -1
4 4    -1
Appendix E Testing $s$-sparsity
In this section our goal is to design an algorithm that given a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ accepts (with probability $\geq\frac{2}{3}$) if $f$ is $s$-sparse and rejects (with probability $\geq\frac{2}{3}$) if $f$ is $\epsilon$-far from any $s$-sparse Boolean function, where $p$ is an odd prime.
The algorithm’s primary concept closely aligns with that presented in [15]. Since we know from Theorem 1.5 that all the coefficients of an $s$-sparse function are more than $1/(p^{2}s)^{\lceil(p-1)/2\rceil}$, so the idea is to partition the Dual group (the set of characters) into $\Theta(s^{2})$ buckets, and for each bucket $B$ estimate the weight $\mathsf{wt}(B)$ of the bucket which is defined as follows:
Definition E.1 (The weight function).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a function. The weight of a bucket $B$ is defined by
$wt(B)\;:=\;\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}.$
We will try to estimate the weight of each bucket within an additive error of $\tau/3$, where $\tau\geq\frac{1}{(p^{2}s)^{p-1}}$. If the function $f$ is indeed $s$-sparse, then clearly the number of buckets with estimated weight more than $32\tau/$ 3 is at most $s$. On the other hand, we will show that if the function $f$ is $\epsilon$-far from any $s$-sparse Boolean function, then with high probability at least $(s+1)$ buckets will have estimated weight more than $32\tau/$ 3.
Thus, the main problem boils down to estimating the weight of a bucket by querying $f$ at a small number of points. Unfortunately, for arbitrary buckets the number of queries to be made can be a lot! To bypass this, the buckets are choosen carefully. The buckets corresponds to the cosets of $H^{\perp}$ where $H$ is a random subspace of $\mathbb{Z}_{p}^{n}$ of dimenstion, $t=\Theta(s^{2})$. For such kind of buckets we show that estimation of the weight can be done using a small number of samples.
E.1 The Algorithm
Algorithm 1 takes a Boolean valued function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ as input. It accepts with probability $\geq\frac{2}{3}$ if $f$ is $s$-sparse, and rejects with probability $\geq\frac{2}{3}$ if $f$ is $\epsilon$-far from any $s$-sparse Boolean valued function.
In lines 4 and 5 we pick a random vector space of $\mathbb{Z}_{p}^{n}$ of dimension $t$, spanned by $v_{1},\ldots,v_{t}$.
For each $b=(b_{1},\ldots,b_{t})$ in $\mathbb{Z}_{p}^{t}$, we define the bucket $C(b)$ as follows:
$C(b):=\{\chi\in\mathbb{Z}_{p}^{n}:\langle\chi,v_{i}\rangle=b_{i}\ \forall i\in\{1,2,\ldots,t\}\}.$
We perform a random shift $u\in\mathbb{Z}_{p}^{t}$.12) and rename the buckets $C(b)$ as $C_{u}(b)$. Observe that the buckets $C_{u}(b)$ gives a partition of $\mathbb{Z}_{p}^{n}$ and each $C_{u}(b)$ corresponds to a coset in $\mathbb{Z}_{p}^{n}/H^{\perp}$. Thus once we have picked the random vector space $H$ in lines 4 and 5 we have a partition of $\mathbb{Z}_{p}^{n}$ (and thus the Dual space of $\mathbb{Z}_{p}^{n}$) into $p^{t}$ buckets where each bucket is of the form $(r+H^{\perp})$ for $r\in\mathbb{Z}_{p}^{n}$.
Now in lines 6 to 9 we estimate the weight of each buckets. The following lemma proved the correctness of the estimation.
Lemma E.2 (Estimating the wt function).
For all $r\in\mathbb{Z}_{p}^{n}$, the value $\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$ calculated in Line 9 satisfy the following:
$\Pr\left(\left|\mathsf{wt}(r+H^{\perp})-\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})\right|\leq\frac{\tau}{3}\right)\geq 1-\frac{1}{10p^{t}}.$
The proof of Lemma E.2 is given in Section E.2. From Lemma E.2 we observe that by union bound (since there are $p^{t}$ buckets) at the end of line 9 with probability at least $\frac{9}{10}$ for all $r\in\mathbb{Z}_{p}^{n}$ the value of $\mathsf{wt}(r+H^{\perp})$ is estimated within additive $\frac{\tau}{3}$. That is, with probability $\frac{9}{10}$
$\forall r\in\mathbb{Z}_{p}^{n},\ \left|\mathsf{wt}(r+H^{\perp})-\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})\right|\leq\frac{\tau}{3}\;.$ (E.1)
Now we can prove the completeness and soundness of our algorithm.
Lemma E.3 (Completeness theorem).
If $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is an $s-$sparse Boolean function, then with probability $\frac{2}{3}$, the test accepts.
Proof E.4.
Let $f=\sum_{\alpha}\hat{f}(\alpha)\chi_{\alpha}$. Since, $f$ is at most $s$-sparse, the support set $A:=\{\alpha\mid\hat{f}(\alpha)\neq 0\}$ has cardinality at most $s$. So clearly, there are at most $s$ buckets (that is at most $s$ different $r$) such that $\mathsf{wt}(r+H^{\perp})$ is non-zero. Also by Theorem 1.5 we know that each non-zero bucket must have weight at least $\tau$.
It thus follows from Equation E.1 that with probability $\frac{9}{10}$, for each bucket with zero weight the estimation is $\leq\frac{\tau}{3}$ and for bucket with non-zero weights the estimation is $\geq\frac{2\tau}{3}$. Thus, in line 12 the algorithm will output YES with probability at least $\frac{9}{10}$.
We now prove the soundness of our algorithm. We wish to show that if $f$ is $\epsilon$-far from any $s$-sparse Boolean function, then the test outputs NO with high probability. For this, we need to prove that in this case there are at least $s+1$ buckets with weight of more than $\tau$. This is proved in the following lemma.
Lemma E.5.
If $f:\mathbb{Z}_{p}^{n}\to\{+1,-1\}$ is $\epsilon$-far from any $s$-sparse Boolean function, where $\epsilon\leq\frac{1}{8(p^{2}s)^{p-1}}$, then with probability $\frac{9}{10}$ there are at least $s+1$ buckets (that is at least $s+1$ vectors $r$ in $\mathbb{Z}_{p}^{n}$) such that weights of these buckets are more that $\tau$.
The proof of Lemma E.5 is given in Section E.3. Using Lemma E.5 we have the soundness of our algorithm.
Lemma E.6.
If $f$ is $\epsilon$-far from any $s$-sparse Boolean function, where $\epsilon\leq\frac{1}{8(p^{2}s)^{p-1}}$, the test rejects with probability $\frac{2}{3}$.
Proof E.7.
Let $f$ be $\epsilon$-far from any $s$-sparse Boolean function. By Lemma E.5 with probability $\frac{9}{10}$ there are at least $(s+1)$ buckets with weight at least $\tau$.
From Equation E.1, with probability $\frac{9}{10}$ the estimations of all the bucket are indeed accurate and thus for each bucket with weight $\geq\tau$ the estimation is $\geq\frac{2\tau}{3}$. And so in line 14 the algorithm will output NO.
Finally we observe that the algorithm makes $O(M)=O(\log(p^{t})\cdot\frac{1}{\tau^{2}})=\mathsf{poly}((p^{2}s)^{p},\frac{1}{\epsilon})$ samples.
E.2 Proof of Lemma E.2
We start by proving the following lemma that proves that on expectation the output of line 9 is $\mathsf{wt}(r+H^{\perp})$.
Lemma E.8.
$\mathbb{E}_{x\in\mathbb{Z}_{p}^{n},z\in H}[\chi_{r}(z)f(x)f(x-z)]=\mathsf{wt}(r+H^{\perp})$.
Proof E.9 (Proof of Lemma E.8).
For the proof of this lemma we need the projection operator defined in Definition B.1. Recall given a subspace $H^{\perp}\subseteq\mathbb{Z}_{p}^{n}$ and a coset $r+H^{\perp}$, we define the projection operator $P_{r+H^{\perp}}f$ on functions $f:\mathbb{Z}_{p}^{n}\to\mathbb{R}$ as follows:
$\widehat{P_{r+H^{\perp}}f}(\alpha)=\begin{cases}\widehat{f}(\alpha)&\alpha\in r+H^{\perp}\\ 0&\text{otherwise}.\end{cases}$
Recall that $wt(r+H^{\perp})=\sum_{\alpha\in r+H^{\perp}}|\hat{f}(\alpha)|^{2}$ which is equal to $\sum_{\alpha\in\mathbb{Z}_{p}^{n}}|\widehat{P_{r+H^{\perp}}f}(\alpha)|^{2}$, by the definition of $\widehat{P_{r+H^{\perp}}f}$. By Parseval, we know that
$\sum_{\alpha\in\mathbb{Z}_{p}^{n}}|\widehat{P_{r+H^{\perp}}f}(\alpha)|^{2}=\mathbb{E}_{w\in\mathbb{Z}_{p}^{n}}[P_{r+H}f(w)\overline{P_{r+H}f(w)}].$
From Lemma B.2 we have $P_{r+H^{\perp}}f(x)=\mathbb{E}_{z\in H}[f(x-z)\chi_{r}(z)]$ and hence $\overline{P_{r+H}f(x)}=\mathbb{E}_{z\in H^{\perp}}[f(x-z)\overline{\chi_{r}(z)}]$. Thus,
$\mathbb{E}_{w\in\mathbb{Z}_{p}^{n}}[P_{r+H}f(w)\overline{P_{r+H}f(w)}]\;=\;\mathbb{E}_{w\in\mathbb{Z}_{p}^{n},z_{1},z_{2}\in H^{\perp}}[\chi_{r}(z_{1})f(w-z_{1})\overline{\chi_{r}(z_{2})}f(w-z_{2})]\;.$
Since, the conjugate of $w_{p}$ ($p$-th root of unity) is $w_{p}^{p-1}$, by definition of $\chi_{r}(\cdot)$, we have
$\chi_{r}(z_{1})\cdot\overline{\chi_{r}(z_{2})}=\chi_{r}(z_{1})\cdot\chi_{r}(-z_{2})=\chi_{r}(z_{1}-z_{2})\;.$
Define, $z:=z_{1}-z_{2}$, and $x:=w-z_{2}$. Then, $x-z=w-z_{1}$. Rewriting the above, we get
$wt(r+H)=\mathbb{E}_{x\in\mathbb{Z}_{p}^{n},z\in H^{\perp}}[\chi_{r}(z)f(x)f(x-z)]\;$
as desired.
We will be using the Hoeffding’s Inequality (Lemma A.22) for our proof of correctness.
Now we are ready to prove Lemma E.2.
Proof E.10 (Proof of Lemma E.2).
Define $\delta:=\frac{1}{10p^{t}}$. Define $Y_{i}:=\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$. Write $\chi_{r}(z_{i})=a_{r}(z_{i})+\iota b_{r}(z_{i})$, where $a_{r}(z_{i})$ and $b_{r}(z_{i})$ are real numbers. Furthermore, since $|\chi_{r}(z_{i})|=1$ (Lemma A.2 (4)), we have that $a_{r}(z_{i}),b_{r}(z_{i})\in[-1,1]$. Define, $Y_{1,i}:=a_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$, and $Y_{2,i}:=b_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$. Since $f$ is Boolean valued, we have that $Y_{1,i},Y_{2,i}\in[-1,1]$. By definition, $Y_{i}=Y_{1,i}+\iota Y_{2,i}$. Consider $Y_{1}:=\frac{\sum_{i\in[M]}Y_{1,i}}{M}$ and $Y_{2}:=\frac{\sum_{i\in[M]}Y_{2,i}}{M}$. Since the pairs $(x_{i},z_{i})$ are all drawn independently, we have that the random variables $\{Y_{1,i}\}_{i=1}^{k}$ are independent, and the random variables $\{Y_{2,i}\}_{i=1}^{k}$ are independent. By Hoeffding’s Inequality (Lemma A.22) we have that for each $i\in[2]$,
$\Pr\left[|Y_{i}-\mathbb{E}[Y_{i}]|\geq\frac{\tau}{\sqrt{2}}\right]$
[…]
where the last inequality follows by an appropriate choice of the constant hidden in the definition $M=O(\log(p^{t})\cdot\frac{1}{\tau^{2}})$ in Algorithm 1. Let $Y=Y_{1}+\iota Y_{2}=\frac{\sum_{i\in[n]}Y_{i}}{M}$. Note that, $(Y-\mathbb{E}[Y])^{2}=(Y_{1}-\mathbb{E}[Y_{1}])^{2}+(Y_{2}-\mathbb{E}[Y_{2}])^{2}$. Trivially, $|Y-\mathbb{E}[Y]|\geq\tau\implies\exists i\in[2]$, such that $|Y_{i}-\mathbb{E}[Y_{i}]|\geq\frac{\tau}{\sqrt{2}}$. Therefore, we have
$\Pr[|Y-\mathbb{E}[Y]|\geq\tau]\leq\Pr[|Y_{1}-\mathbb{E}[Y_{1}]|\geq\frac{\tau}{\sqrt{2}}]+\Pr[|Y_{2}-\mathbb{E}[Y_{2}]|\geq\frac{\tau}{\sqrt{2}}]\;\leq\;\delta\;.$
From Lemma E.8 we have that $\mathbb{E}[Y]=\mathsf{wt}(r+H^{\perp})$. Hence we have our lemma.
E.3 Proof of Lemma E.5
Before we go on to prove Lemma E.5 it is important to understand some structure of the partition of the dual set we obtain in Lines 5 and 9. Recall that after we select the subspace $H$ at random the dual set is partitioned into buckets $C(b)$, where,
$C(b)=\{\chi\in\mathbb{Z}_{p}^{n}:\langle\chi,v_{i}\rangle=b_{i}\ \forall i\in\{1,2,\ldots,t\}\},$
where $v_{i}$ are the rows of the matrix $A$.
Consider the following random variables:
Definition E.11.
The indicator random variable $I_{\chi\to b}$ for the event $\chi\in C(b)$ is defined by
$I_{\chi\to b}=\begin{cases}1&\mbox{ If }\chi\in C(b)\\ 0&\text{otherwise}.\end{cases}$
Definition E.12 (Random shift).
Let us pick a non-zero element $u\in\mathbb{Z}_{p}^{t}$. For each $b$, let us define $C(b+u)$ as $C_{u}(b)$. This is known as a random shift. It is easy to see that $C_{u}(b)$ is also a coset and is isomorphic to $C(b)$.
Recall the definition of covariance of two random variables (Definition A.19). We need the following theorem.
Theorem E.13.
Let us consider the partition in buckets $C(b)$. Then the following are true:
  1. 1.
    For all $\chi\in\mathbb{Z}_{p}^{n}$ and for all $b\in\mathbb{Z}_{p}^{t}$, $\Pr_{u,A}[\chi\in C_{u}(b)]=\frac{1}{p^{t}}$.
  2. 2.
    Let $S\subseteq\mathbb{Z}_{p}^{n}$ be such that $|S|\leq s+1$. Then, if $t\geq 2\log_{p}s+\log_{p}(\frac{1}{\delta})$, all $\chi\in S$ belong to different buckets except with probability at most $\delta$.
  3. 3.
    For all $b\in\mathbb{Z}_{p}^{t}$ and for all distinct $\chi,\chi^{\prime}\in\widehat{\mathbb{Z}_{p}^{n}}$, $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=0$.
Proof E.14.
  1. 1.
    The event $\chi\in C_{u}(b)$ is equivalent to the event $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$. For each $i\in[t]$, $\Pr[\langle\chi,v_{i}\rangle=b_{i}+u_{i}]=\Pr[u_{i}=\langle\chi,v_{i}\rangle-b_{i}]=1/p$. Since, $u_{1},u_{2},\ldots,u_{t}$ and $b_{1},\ldots,b_{t}$ are all independent, the statement follows.
  2. 2.
    Let $\chi$ and $\chi^{\prime}$ be two distinct characters in $\widehat{\mathbb{Z}_{p}^{n}}$. We have that
    $\Pr\left[\chi,\chi^{\prime}\text{ belong to the same bucket}\right]=\Pr\left[\forall i\in[t],\langle\chi-\chi^{\prime},v_{i}\rangle=0\right]=\frac{1}{p^{t}},$
    where the last equality holds because $\chi$ and $\chi^{\prime}$ are distinct, and $v_{1},\ldots,v_{t}$ are independent.
    Since $|S|\leq s+1$, therefore the number of ways in which two distinct $\chi,\chi^{\prime}$ can be chosen from $S$ is $\binom{s+1}{2}\leq s^{2}$. Therefore, probability that all $\chi\in S$ belong to different buckets is given by
    Pr[all $\Pr[\text{all }\chi\in S\text{ belong to different buckets}]$ $\geq 1-\binom{s+1}{2}\frac{1}{p^{t}}$
    $\geq 1-s^{2}\frac{1}{p^{t}}$
    $\geq 1-s^{2}\frac{1}{p^{2\log_{p}s+\log_{p}(\frac{1}{\delta})}}$
    $=1-s^{2}\frac{1}{\frac{s^{2}}{\delta}}$
    $=1-\delta.$
  3. 3.
    For all $b$, $\Pr_{u}[b+u\neq 0]=\frac{p^{t}-1}{p^{t}}$. Now, let us look at the following cases to prove the theorem.
    Case 1: $\chi^{\prime}$ is a scalar multiple of $\chi$, and both of them are nonzero.
    In this case,
    $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if
    1. (a)
      $b+u=0^{t}$, i.e., $u=-b$, and
    2. (b)
      $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$.
    The probability of event (a) is $\frac{1}{p^{t}}$. Now, conditioned on event (a). Since the $v_{1},\ldots,v_{t}$ and $u$ are all independent random variables, the conditional probability of event (b) is $\frac{1}{p^{t}}$. We thus have that $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{t}}\cdot\frac{1}{p^{t}}=\frac{1}{p^{2t}}$.
    Now, $\mathbb{E}[I_{\chi\to b+u}]=\Pr[\chi\in C_{u}(b)]=\frac{1}{p^{t}}$ by part (1). Similarly, $\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{t}}$. We thus have that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=\frac{1}{p^{2t}}-\frac{1}{p^{t}}\cdot\frac{1}{p^{t}}=0$.
    Case 2: $\chi,\chi^{\prime}$ are both nonzero, distinct and linearly independent.
    In this case,
    $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$ and $\langle\chi^{\prime},v_{i}\rangle=b_{i}+u_{i}$. Fix $i\in[t]$ and $u_{i}$. Since $\chi$ and $\chi^{\prime}$ are linearly independent, the simultaneous solutions of the linear equations $\langle\chi,v_{i}\rangle=b_{i}+u_{i}$ and $\langle\chi^{\prime},v_{i}\rangle=b_{i}+u_{i}$ span an affine subspace of $\mathbb{Z}_{p}^{n}$ of co-dimension $2$. Thus, for each $i$, the probability that $v_{i}$ belongs to this affine subspace is $\frac{1}{p^{2}}$. Since $u_{1},\ldots,u_{t},v_{1},\ldots,v_{t}$ are all independent, we have that $\mathbb{E}[I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{2t}}$.
    Now, for the same reason as in the previos case, $\mathbb{E}[I_{\chi\to b+u}]=\mathbb{E}[I_{\chi^{\prime}\to b+u}]=\frac{1}{p^{t}}$. we thus have that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=\frac{1}{p^{2t}}-\frac{1}{p^{t}}\cdot\frac{1}{p^{t}}=0$.
    So, analyzing all possible cases, we observe that $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=0$.
We will also need the following application of Chebyshev’s inequality.
Lemma E.15.
Let $X=\sum_{i\in[n]}X_{i}$, where the $X_{i}$’s are real random variables satisfying – (i) $0\leq X_{i}\leq\tau$, and (ii) $\mathsf{cov}(X_{i},X_{j})=0$, for $i\neq j$. Further, assume that $\mathbb{E}[X]>0$. Then for any $\epsilon^{\prime}>0$, we have $\operatorname{Var}[X]\leq\tau\mathbb{E}[X]$ and
$\Pr[X\leq(1-\epsilon^{\prime})\mathbb{E}[X]]\;\leq\;\frac{\tau}{\epsilon^{\prime 2}\mathbb{E}[X]}.$
Proof E.16.
$\operatorname{Var}[X]$ $=\sum_{i}\operatorname{Var}[X_{i}]+\operatorname{Cov}_{i,j:i\neq j}[X_{i}X_{j}]$
$=\sum_{i}\mathbb{E}[X_{i}^{2}],\text{ since }\operatorname{Cov}_{i,j:i\neq j}[X_{i}X_{j}]=0$
$\leq\sum_{i}\tau\mathbb{E}[X_{i}]$
$=\tau\mathbb{E}[X].$
Therefore, by Chebyshev’s inequality,
$\Pr[X\leq(1-\epsilon^{\prime})\mathbb{E}[X]]$ $=\Pr[X-\mathbb{E}[X]\leq-\epsilon^{\prime}\mathbb{E}[X]]$
$\leq\Pr[|X-\mathbb{E}[X]|\geq\epsilon^{\prime}\mathbb{E}[X]]$
$\leq\frac{\operatorname{Var}[X]}{\epsilon^{\prime 2}\mathbb{E}[X]^{2}}\leq\frac{\tau\mathbb{E}[X]}{\epsilon^{\prime 2}\mathbb{E}[X]^{2}}=\frac{\tau}{\epsilon^{\prime 2}\mathbb{E}[X]}.$
Finally we are ready to prove Lemma E.5.
Proof E.17 (Proof of Lemma E.5).
To prove this lemma, let us divide the Fourier coefficients of $f$ in two sets:
$A_{1}\;:=\;\{\chi:|\widehat{f}(\chi)|^{2}\geq\tau\},\ A_{2}\;:=\;\{\chi:|\widehat{f}(\chi)|^{2}<\tau\}\;.$
Since $f$ is $\epsilon$-far from any $s$-sparse Boolean function with $\epsilon\leq\frac{1}{8(p^{2}s)^{p-1}}$, so by Theorem 1.4, it is $\frac{\epsilon}{\sqrt{2}}$-far from any $s$-sparse function. We will prove the lemma in two cases.
Case 1: $|A_{1}|\geq s+1$.
In this case from Theorem E.13 (part (2)) implies that with probability at least $\frac{19}{20}$, at least $s+1$ buckets will contain elements from $A_{1}$. Therefore, the weights of those buckets will be at least $\tau$, and hence it follows.
Case 2: $|A_{1}|\leq s$ $\mathsf{wt}(A_{2})\geq\epsilon^{2}/2$.
Let us define the random variables $Y_{b}=\mathsf{wt}(C_{u}(b)\cap A_{2})$ for each $b$. Then $Y_{b}=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$, where $I_{\chi\to b+u}$ is the indicator random variable which takes the value $1$ if $\chi\in C_{u}(b)$ and 00 otherwise.
To show at least $(s+1)$ of the buckets have weight more than $\tau$ we will first prove that on expectation the weight of a bucket is significantly more than $\tau$. Then we will use Chebyshev’s inequality to prove that the probability that a bucket has less than $\tau$ is small and finally, we will use Markov’s inequality to upper bound the number of small buckets - or in other words prove that at least $(s+1)$ buckets have weight more than $\tau$.
We start by proving that on expectation the weights of the buckets are large. Note that,
$\mathbb{E}[Y_{b}]=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\Pr[\chi\in C_{u}(b)]=\frac{1}{p^{t}}\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\geq\frac{\epsilon^{2}}{2p^{t}},$ (E.2)
where the second equality follows from the fact that $\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\geq(\frac{\epsilon}{\sqrt{2}})^{2}$ and from Theorem E.13 (part (1)) where we show that $\Pr[\chi\in C_{u}(b)]=\frac{1}{p^{t}}$.
Before we apply Chebyshev’s inequality to upper bound the probability that a bucket has small weight we upper bound the variance of $Y_{b}$. We can write $Y_{b}=\sum_{\chi\in A_{2}}Y_{\chi,b}$, where $Y_{\chi,b}=|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$. Note that $0\leq Y_{b}\leq\tau$, $\mathbb{E}[Y_{b}]>0$ (by Equation E.2) and $\mathsf{cov}(Y_{\chi,b},Y_{\chi^{\prime},b})=0$ (by Theorem E.13, where $\chi,\chi^{\prime}\in A_{2}$). So, by Lemma E.15,
$\operatorname{Var}[Y_{b}]\leq\tau\mathbb{E}[Y_{b}].$
Therefore, by Chebychev’s inequality we have that, for any $\epsilon^{\prime}$,
$\Pr[Y_{b}\leq(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]]$ $\leq\frac{\operatorname{Var}[Y_{b}]}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]^{2}}\leq\frac{\tau\mathbb{E}[Y_{b}]}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]^{2}}=\frac{\tau}{\epsilon^{\prime 2}\mathbb{E}[Y_{b}]}\leq\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\cdot\epsilon^{2}}.$
Since the total number of buckets is $p^{t}$, therefore the expected number of buckets with weight $\leq(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $p^{t}\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}$. Thus by Markov’s inequality, with probability at least $\frac{9}{10}$, the number of buckets with weight more that $(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}$.
Now, set $\epsilon^{\prime}:=\frac{19}{20}$. Thus,
$(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]\geq\frac{1}{20}\cdot\frac{\epsilon^{2}}{2p^{t}}=\frac{\epsilon^{2}}{40p^{t}}\geq\tau.$
It thus follows that with probability at least $\frac{9}{10}$, the number of buckets with weight more that $(1-\epsilon^{\prime})\mathbb{E}[Y_{b}]$ is at most $10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}$.
We now finish by showing that $10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}\leq p^{t}-(s+1)$.
$10\cdot\frac{\tau\cdot 2p^{t}}{\epsilon^{\prime 2}\epsilon^{2}}\cdot p^{t}\leq p^{t}-(s+1)$
$\Longleftarrow\tau\leq\frac{\epsilon^{2}}{p^{t}}\cdot\biggl{(}1-\frac{s+1}{p^{t}}\biggr{)}\cdot\left(\frac{19}{20}\right)^{2}\cdot\frac{1}{20}$
$\Longleftarrow\tau\leq\frac{\epsilon^{2}}{p^{t}}\cdot\frac{1}{2}\cdot\frac{1}{20}=\frac{\epsilon^{2}}{40p^{t}}.$
where the last implication follows from the fact that $(1-\frac{s+1}{p^{t}})\geq\frac{1}{2}\cdot\left(\frac{20}{19}\right)^{2}$ for all $s$, since $t=\lceil\log_{p}(20s^{2})\rceil+1$. We have set $\tau\leq\frac{\epsilon^{2}}{40p^{t}}$ in the algorithm. Hence, the number of heavy buckets is at least $(s+1)$.
Appendix F Proof of Theorem F.8
In this section we will prove the generalized version of Theorem 1.3 and Theorem 1.4 for $\mathcal{G}=\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$, where $p_{1},\ldots,p_{t}$ are distinct primes. Let $\omega_{p_{i}}$ be a primitive $p_{i}^{th}$ root of unity, $i\in[t]$. For any multivariate polynomial $\sum_{i_{1}=0}^{d_{1}}\cdots\sum_{i_{t}=0}^{d_{t}}g_{i_{1},\ldots,i_{t}}x_{1}^{i_{1}}\cdots x_{t}^{i_{t}}=g(x_{1},\ldots,x_{t})\in\mathbb{C}[x_{1},\ldots,x_{t}]$, let $|g|_{1}\;:=\;\sum_{i_{1}=0}^{d_{1}}\cdots\sum_{i_{t}=0}^{d_{t}}|g_{i_{1},\ldots,i_{t}}|$.
The following definition helps us to understand what we mean by $\mu$-close to $(m_{1},\ldots,m_{t})$-granular.
Definition F.1.
($\mu$-close to $(m_{1},\ldots,m_{t})$-granular) A complex number $v$ is said to be $\mu$-close to $(m_{1},\ldots,m_{t})$-granular with respect to $\mathbb{Z}_{p_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}$ if $|v-\frac{g(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|\leq\mu$, for some $g(X_{1},\ldots,X_{t})\in\mathbb{Z}[X_{1},\ldots,X_{t}]$, which is a non-zero multivariate polynomial of degree at most $(p_{1}-1)\cdots(p_{t}-1)$, and $\omega_{p_{i}}$ is a $p_{i}^{th}$ root of unity, $i\in[t]$.
Now let us state the generalized form of Theorem 1.3 for $\mathcal{G}$.
Theorem F.2.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function. Then, if $f$ is $\mu$-close to $s$-sparse in $\ell_{2}$, for each character $\chi\in B$, $\widehat{f}(\chi)$ is $\frac{\mu}{\sqrt{s}}$-close to $(\lceil\log_{p_{1}}(s)^{\frac{1}{t}}\rceil+1,\ldots,\lceil\log_{p_{t}}(s)^{\frac{1}{t}}\rceil+1)$-granular.
The following theorem is a corollary of Theorem F.2 which gives us a lower bound on the Fourier coefficients of $f:\mathcal{G}\to\{-1,+1\}$.
Theorem F.3 (Theorem F.8, restated).
Let $p_{1},\ldots,p_{t}$ be distinct primes, and let $f:\mathbb{Z}_{p_{1}}^{n_{1}}\times\mathbb{Z}_{p_{2}}^{n_{2}}\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}\to\{-1,+1\}$. Then, for each $\chi\in\operatorname{supp}(f)$, we have
$\left|\widehat{f}(\chi)\right|\geq\frac{1}{(m^{2}s_{f})^{\lceil\varphi(m)/2\rceil}},$
where $m:=p_{1}\dots p_{t}\text{ and }\varphi(m)=(p_{1}-1)(p_{2}-1)\cdots(p_{t}-1)$ and $s_{f}$ is the sparsity of the function $f$.
The following theorem is the generalized form of Theorem 1.4 for $\mathcal{G}$.
Theorem F.4.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function such that $f$ is $\mu\leq\frac{1}{8(m^{2}s)^{\varphi(m)}}$ close to $s$-sparse in $\ell_{2}$ norm, where $m=p_{1}\cdots p_{t}$ and $\varphi(m)=(p_{1}-1)\cdots(p_{t}-1)$. Then, there exists an $s$-sparse Boolean valued function $F:\mathcal{G}\to\{-1,+1\}$ which is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$.
To prove Theorem F.2, we need the following lemma.
Lemma F.5.
$P_{r_{1}+H_{1},\ldots,r_{t}+H_{t}}f(x_{1},\ldots,x_{t})=\frac{1}{|H_{1}^{\perp}\cdots H_{t}^{\perp}|}\sum_{(z_{1},\ldots,z_{t})\in H_{1}^{\perp}\cdots H_{t}^{\perp}}f(x_{1}-z_{1},\ldots,x_{t}-z_{t})\chi((z_{1},\cdots,z_{t})),$
where $\chi$ is a character of $\mathcal{G}$ and for each $i$, $H_{i}$ is a subspace (hence a subgroup) of $\mathbb{Z}_{p_{i}}^{n_{i}}$ and $x_{i},z_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$.
The following lemma plays a crucial role in proving Lemma F.5.
Claim 4 (Restated Lemma A.10).
$\sum_{h\in H}\chi_{h}=|H|\cdot 1_{H^{\perp}}$, where $H=H_{1}\times\cdots\times H_{t}$ is a subgroup of $\mathcal{G}$, and $h=(h_{1},\ldots,h_{t}),\ h_{i}\in H_{i}\ \forall i$. That is,
$\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\chi_{h_{1}}\cdots\chi_{h_{t}}=|H_{1}|\cdots|H_{t}|\cdot 1_{H_{1}^{\perp}}\cdots 1_{H_{t}^{\perp}}.$
Here, for each $i$, $H_{i}$ is a subgroup of $\mathbb{Z}_{p_{i}}^{n_{i}}$ and hence a subspace $\mathbb{Z}_{p_{i}}^{n_{i}}$, as $\mathbb{Z}_{p_{i}}^{n_{i}}$ is a vector space.
Proof F.6 (Proof of Lemma F.5).
$P_{r_{1}+H_{1},\ldots,r_{t}+H_{t}}f(x_{1},\ldots,x_{t})$
$=\sum_{\alpha_{1}\in r_{1}+H_{1},\ldots,\alpha_{t}\in r_{t}+H_{t}}\hat{f}(\alpha_{1},\ldots,\alpha_{t})\chi^{(p_{1})}_{\alpha_{1}}(x_{1})\cdots\chi^{(p_{t})}_{\alpha_{t}}(x_{t})$
$=\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\hat{f}(r_{1}+\beta_{1},\ldots r_{t}+\beta_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}+\beta_{i}}(x_{i}),\text{ where }\alpha_{i}=r_{i}+\beta_{i},i\in\{1\ldots,t\}\text{ and }\beta_{i}\in H_{i}\forall i$
$=\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\hat{f}(r_{1}+\beta_{1},\ldots,r_{t}+\beta_{t})\prod_{i=1}^{t}\bigg{(}\chi^{(p_{i})}_{r_{i}}(x_{i})\chi^{(p_{i})}_{\beta_{i}}(x_{i})\bigg{)}$
$=\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\hat{f}(r_{1}+\beta_{1},\ldots,r_{t}+\beta_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i})$
$=\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i})\frac{1}{|\mathcal{G}|}\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\overline{\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}+\beta_{i}}(y_{i})}$
$=\frac{1}{|\mathcal{G}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i})\prod_{i=1}^{t}\overline{\chi^{(p_{i})}_{r_{i}+\beta_{i}}(y_{i})}$
$=\frac{1}{|\mathcal{G}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(-y_{i})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i}-y_{i})$
$=\frac{1}{|\mathcal{G}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\bigg{(}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(-y_{i})\bigg{)}|H_{1}|\cdots|H_{t}|1_{H_{1}^{\perp}}(x_{1}-y_{1})\cdots 1_{H_{t}^{\perp}}(x_{t}-y_{t})$
$=\frac{1}{|H_{1}^{\perp}|\cdots|H_{t}^{\perp}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{x_{1}-y_{1}\in H_{1}^{\perp},\ldots x_{t}-y_{t}\in H_{t}^{\perp}}f(y_{1},\ldots,y_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(-y_{i})$
$=\frac{1}{|H_{1}^{\perp}|\cdots|H_{t}^{\perp}|}\sum_{x_{1}-y_{1}\in H_{1}^{\perp},\ldots,x_{t}-y_{t}\in H_{t}^{\perp}}f(y_{1},\ldots,y_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i}-y_{i})$
$=\frac{1}{|H_{1}^{\perp}|\cdots|H_{t}^{\perp}|}\sum_{z_{1}\in H_{1}^{\perp},\ldots z_{t}\in H_{t}^{\perp}}f(x_{1}-z_{1},\ldots,x_{t}-z_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(z_{i}),$
where $x_{i}-y_{i}=z_{i}$ for all $i\in[t]$.
Now we prove Theorem F.2.
Proof F.7 (Proof of Theorem F.2).
Fix a $\chi^{(\mathcal{G})}:=\chi^{(p_{1})}\dots\chi^{(p_{t})}\in B$, where $\chi^{(p_{i})}$ is a character of $\mathbb{Z}^{n_{i}}_{p_{i}}$ for all $i\in[t]$, and $B$ is $B$ is the set of character functions corresponding to the $s$ largest coefficients. Set for each $i\in[t]$, $m_{i}:=\lceil\log_{p_{i}}(s)^{\frac{1}{t}}\rceil+1$.
Let $A_{i}^{\perp}\subset\mathbb{Z}^{n_{i}}_{p_{i}}$, be the set of solutions to $A_{i}\chi=0$, for all $i\in[t]$. Let $H_{i}$ be the coset of $A_{i}^{\perp}$ that are solutions to system of linear equations $A_{i}\chi=b_{i}$. From the definition of projection operator, we have
$P_{H_{1},\ldots,H_{t}}f(x_{1},\ldots,x_{t})=\sum_{\chi^{(p_{1})}\in H_{1},\ldots,\chi^{(p_{t})}\in H_{t}}\widehat{f}(\chi^{(p_{1})},\ldots,\chi^{(p_{t})})\chi^{(p_{1})}(x_{1})\cdots\chi^{(p_{t})}(x_{t}),$ (F.1)
where $\chi_{r_{j}}^{(p_{i})}$ is a character of $\mathbb{Z}_{p_{i}}^{n_{i}}$ (follows from Definition A.1).
We need to show that for each $\chi_{r_{j}}=(\chi_{r_{j}}^{(p_{1})}\cdots\chi_{r_{j}}^{(p_{t})})\in B$, where $\chi_{r_{j}}^{(p_{i})}$ is a character of $\mathbb{Z}_{p_{i}}^{n_{i}}$, there exists $m_{i}\times n_{i}$ matrices $A_{i}$ and a $m_{i}$-dimensional vector $b_{i}$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A_{1}\chi^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi^{(p_{t})}\,=\,b_{t}\right\}=\{\chi_{r_{j}}\},$ (F.2)
and, the $\ell_{2}$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, where $H=H_{1}\times\cdots\times H_{t}$, that is,
$\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\leq\frac{\mu^{2}}{s},$
where $B$ is the set of character functions corresponding to the $s$ largest coefficients and $S$ is the set of rest of the characters, and $H_{i}$ is the solution set of the system of linear equations $A_{i}\chi=b_{i}$ for all $i\in[t]$. Let $E_{1}$ and $E_{2}$ be the first and the second events respectively. So we need to prove that both these events occur simultaneously with non-zero probability.
To show the existence of $A_{i}$ and $b_{i}$ for all $i\in[t]$, given $\chi_{r_{j}}\in B$, for which Equation F.2 is true we will need to consider two cases:
Case 1: $\chi_{r_{j}}=0$
. Observe that for all $A_{i}$, $i\in[t]$, we have $A_{i}\chi_{r_{j}}^{(p_{i})}=0$. If entries of $A_{i}$ are sampled independently and uniformly at random from $\mathbb{Z}_{p_{i}}$, for each $i\in[t]$. Then, for all $\chi_{r_{q}}\in B\setminus\{\chi_{r_{j}}\}$, we have
$\Pr_{A_{1},\ldots,A_{t}}\left[A_{1}\chi_{r_{q}}^{(p_{1})}=0,\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}=0\right]=\frac{1}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}\leq\frac{1}{s}.$
Note that $\Pr_{A_{1},\ldots,A_{t}}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when $A_{i}$ are the random matrices generated by the above-mentioned procedure. Using union bound, we get
$\Pr_{A_{1},\ldots,A_{t}}\left[\exists q\neq j\text{ such that }A_{1}\chi_{r_{q}}^{(p_{1})}=0,\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}=0\right]\leq\frac{s-1}{s}<1.$
Therefore, if $\chi_{r_{j}}^{(p_{i})}=0$ for all $i\in[t]$, then there exists $m_{i}\times n_{i}$ matrices $A_{i}$ such that
[…]
for all $i\in[t]$.
Case 2: $\chi_{r_{j}}\neq 0$
. Fix nonzero vectors $b_{i}\in\mathbb{Z}^{n_{i}}_{p_{i}}\setminus\{(0,\dots,0)\}$ for all $i\in[t]$. Again, let $A_{i}$ be random $m_{i}\times n_{i}$ matrices with entries of $A_{i}$ being sampled independently and uniformly at random from $\mathbb{Z}_{p_{i}}$. Observe that $\Pr_{A_{1},\ldots,A_{t}}[A_{1}\chi_{r_{j}}^{(p_{1})}=b_{1},\ldots,A_{t}\chi_{r_{j}}^{(p_{t})}=b_{t}]=\frac{1}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}$. Let $\chi_{r_{q}}\in B\setminus\{\chi_{r_{j}}\}$. If $\chi_{r_{q}}^{(p_{i})}=\alpha_{i}\chi_{r_{j}}^{(p_{i})}$, where $\alpha_{i}\in\mathbb{Z}_{p_{i}}\setminus\{1\}$ for at least one $i\in[t]$, then
$\Pr_{A_{1},\ldots,A_{t}}[A_{1}\chi_{r_{q}}^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}\,=\,b_{t}\,\mid\,A_{1}\chi_{r_{j}}^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi_{r_{j}}^{(p_{t})}\,=\,b_{t}]=0.$
Now we will consider the case when $\chi_{r_{q}}^{(p_{i})}$ are not scalar multiples of $\chi_{r_{j}}^{(p_{i})}$ for all $i\in[t]$. Observe that
$\Pr_{A_{1},\ldots,A_{t}}[A_{1}\chi_{r_{q}}^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}\,=\,b_{t}\,\mid\,A_{1}\chi_{r_{j}}^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi_{r_{j}}^{(p_{t})}\,=\,b_{t}]$ (F.3)
$=\frac{\Pr_{A_{1},b_{1}}[A_{1}\chi_{r_{q}}^{(p_{1})}\,=\,b_{1},A_{1}\chi_{r_{j}}^{(p_{1})}\,=\,b_{1}]}{\Pr_{A_{1}}[A_{1}\chi_{r_{j}}^{(p_{1})}\,=\,b_{1}]}\cdots\frac{\Pr_{A_{t},b_{t}}[A_{t}\chi_{r_{q}}^{(p_{t})}\,=\,b_{t},A_{t}\chi_{r_{j}}^{(p_{t})}\,=\,b_{t}]}{\Pr_{A_{t}}[A_{t}\chi_{r_{j}}^{(p_{t})}\,=\,b_{t}]}$
$=\prod_{i=1}^{t}\frac{\Pr_{A_{i}}\left[\bigcap\limits_{\ell\in[m_{i}]}\left(A_{i\ell}\cdot\chi_{r_{q}}^{(p_{i})}=b_{i\ell},\,A_{i\ell}\cdot\chi_{r_{j}}^{(p_{i})}=b_{i\ell}\right)\right]}{\Pr_{A_{i}}[A_{i}\chi_{r_{j}}^{(p_{i})}\,=\,b_{i}]}$
$=\prod_{i=1}^{t}\frac{\prod\limits_{\ell\in[m_{i}]}\Pr_{A_{i}}\left[A_{i\ell}\cdot\chi_{r_{q}}^{(p_{i})}=b_{i\ell},\,A_{i\ell}\cdot\chi_{r_{j}}^{(p_{i})}=b_{i\ell}\right]}{\Pr_{A_{i}}[A_{i}\chi_{r_{j}}^{(p_{i})}\,=\,b_{i}]}.$ (F.4)
Note that $A_{i\ell}$ and $b_{i\ell}$ denote the $\ell$-th row of $A_{i}$ and $\ell$-th coordinate of $b_{i}$ respectively, and $[m_{i}]$ denotes the set $\{1,\dots,m_{i}\}$ for all $i\in[t]$. Observe that the last equality in the above equation follows from the fact that rows of $A_{i}$ are independent and uniformly random samples of $\mathbb{Z}^{n_{i}}_{p_{i}}$. As $\chi_{r_{q}}$ and $\chi_{r_{j}}$ are not scalar multiple of each other therefore
$\Pr_{A_{i}}\left[A_{i\ell}\cdot\chi_{r_{q}}=b_{i\ell},\,A_{i\ell}\cdot\chi_{r_{j}}=b_{i\ell}\right]=\frac{1}{p_{i}^{2}},$
for all $i\in[t]$. Using the above equation and the fact that $\Pr_{A_{i}}[A_{i}\chi_{r_{j}}=b_{i}]=\frac{1}{p_{i}^{m_{i}}}$, for each $i\in[t]$ in Equation (F.3) we get
$\Pr_{A_{1},\ldots,A_{t}}[A_{1}\chi_{r_{q}}^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}\,=\,b_{t}\,\mid\,A_{1}\chi_{r_{j}}^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi_{r_{j}}^{(p_{t})}\,=\,b_{t}]$
$=\prod_{i=1}^{t}\frac{\prod\limits_{\ell\in[m_{i}]}\Pr_{A_{i}}\left[A_{i\ell}\cdot\chi_{r_{q}}^{(p_{i})}=b_{i\ell},\,A_{i\ell}\cdot\chi_{r_{j}}^{(p_{i})}=b_{i\ell}\right]}{\Pr_{A_{i}}[A_{i}\chi_{r_{j}}^{(p_{i})}\,=\,b_{i}]}$
$=\prod_{i=1}^{t}\frac{\bigg{(}\frac{1}{p_{i}^{2}}\bigg{)}^{m_{i}}}{\frac{1}{p_{i}^{m_{i}}}}$
$=\frac{1}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}.$
Using the conditional variant of the union bound we get
$\Pr_{A_{1},\ldots,A_{t}}\left[\exists q\neq j\text{ such that }A_{1}\chi_{r_{q}}^{(p_{1})}=0,\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}=0\right]\leq\frac{s-1}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}<1.$
From the above two cases we can now say that given any $\chi_{r_{j}}\in B$, there exists $m_{i}\times n_{i}$ matrices $A_{i}$ and a $m_{i}$-dimensional vector $b_{i}$ for each $i\in[t]$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A_{1}\chi^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi^{(p_{t})}\,=\,b_{t}\right\}=\{\chi_{r_{j}}\}.$
Now, let us look at the coefficients corresponding to the characters in $S$, in order to prove that the $\ell$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, where $H=H_{1}\times\cdots\times H_{t}$, $H_{i}$ being the solution set of the system of linear equations $A_{i}\chi=b_{i}$, $i\in[t]$. Let us look at the expected value of $\sum_{\chi\in S\cap H}\widehat{f}(\chi)^{2}$, given $A_{i}\chi_{r_{j}}=b_{i}$, $i\in[t]$.
$\mathbb{E}_{A_{1},\ldots,A_{t},b_{1},\ldots,b_{t}}\biggl{[}\sum_{\chi\in S\cap H}\widehat{f}(\chi)^{2}\mid A_{1}\chi_{r_{j}}^{(p_{1})}=b_{1}\cdots A_{t}\chi_{r_{j}}^{(p_{t})}=b_{t}\biggr{]}$
$=\sum_{\chi\in S}\Pr_{A_{1},\ldots,A_{t},b_{1},\ldots,b_{t}}[\chi\in H\mid A_{1}\chi_{r_{j}}^{(p_{1})}=b_{1}\cdots A_{t}\chi_{r_{j}}^{(p_{t})}=b_{t}]\widehat{f}(\chi)^{2}$
$\leq\frac{\mu^{2}}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}$
$\leq\frac{\mu^{2}}{p_{1}\cdots p_{t}s},$ (F.5)
since $m_{i}=\lceil\log_{p_{i}}(s)^{\frac{1}{t}}\rceil+1$.
Therefore, by Markov’s inequality, we have,
$\Pr_{A_{1},\ldots,A_{t},b_{1},\ldots,b_{t}}[\sum_{\chi\in S\cap H}\widehat{f}(\chi)^{2}\geq\frac{\mu^{2}}{s}\mid A_{1}\chi_{r_{j}}^{(p_{1})}=b_{1}\cdots A_{t}\chi_{r_{j}}^{(p_{t})}=b_{t}]\leq\frac{\frac{\mu^{2}}{p_{1}\cdots p_{t}s}}{\frac{\mu^{2}}{s}}=\frac{1}{p_{1}\cdots p_{t}}<1.$ (F.6)
So, if we apply union bound to Equation F.2 and Equation F.6, we get that the events $E_{1}$ and $E_{2}$ occur simultaneously with non-zero probability.
Therefore, we have,
$P_{H}f(x)=\widehat{f}(\chi_{r_{j}})\chi_{r_{j}}(x)+\sum_{\chi\in S\cap H}\widehat{f}(\chi)\chi(x),$
where $\sum_{\chi\in S\cap H}\widehat{f}(\chi)^{2}\leq\frac{\mu^{2}}{s}$.
Now, by Lemma B.2, we have,
$P_{H}f(x)=\mathbb{E}_{z\in r(A_{1})\times\cdots\times r(A_{t})}[f(x-z)\chi_{r}(z)],$
which is of the form $\frac{g(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}$, where $g(X_{1},\ldots,X_{t})\in\mathbb{Z}[X_{1},\ldots,X_{t}]$ is a nonzero multivariate polynomial with degree $\leq(p_{1}-1)\cdots(p_{t}-1)$, and $r(A_{i})$ is the row-space of $A_{i}$, $i\in[t]$. So $P_{H}f(x)$ is $(m_{1},\ldots,m_{t})$-granular.
Also, let us look at the function $G_{1}(x)=\sum_{\chi\in S\cap H}\widehat{f}(\chi)\chi(x)$. We know, by F.7, that, $\mathbb{E}_{x}G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$. So, $G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$ for some $x\in\mathcal{G}$. For this particular $x$,
$P_{H}f(x)=\widehat{f}(\chi_{r_{j}})\chi_{r_{j}}(x)+G_{1}(x).$
Therefore,
$|\widehat{f}(\chi_{r_{j}})|=|P_{H}f(x)-G_{1}(x)|,$
which implies that $\widehat{f}(\chi_{r_{i}})$ is $\frac{\mu}{\sqrt{s}}$ close to $(m_{1},\ldots,m_{t})$-granular, since $P_{H}f(x)$ is $(m_{1},\ldots,m_{t})$-granular and $|G_{1}(x)|\leq\frac{\mu}{\sqrt{s}}$.
Now let us prove the generalized lower bound result over $\mathcal{G}$.
Theorem F.8 (Generalized Fourier coefficient lower bound).
Let $p_{1},\ldots,p_{t}$ be distinct primes, and let $f:\mathbb{Z}_{p_{1}}^{n_{1}}\times\mathbb{Z}_{p_{2}}^{n_{2}}\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}\to\{-1,+1\}$. Then, for each $\chi\in\operatorname{supp}(f)$, we have
$\left|\widehat{f}(\chi)\right|\geq\frac{1}{(m^{2}s_{f})^{\lceil\varphi(m)/2\rceil}},$
where $m:=p_{1}\dots p_{t}\text{ and }\varphi(m)=(p_{1}-1)(p_{2}-1)\cdots(p_{t}-1)$.
Proof F.9 (Proof of Theorem F.8).
If we put $\mu=0$ in Theorem F.2, we get that there exists a $g\in\mathbb{Z}[X_{1},\ldots,X_{t}]$, such that $|\widehat{f}(\chi)|\geq|g(\omega_{p_{1}},\ldots,\omega_{p_{t}})/p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}|$, where $m_{i}=\lceil\log_{p_{i}}s_{f}^{\frac{1}{t}}\rceil+1$ for all $i\in[t]$, $s_{f}$ being the sparsity of the Boolean valued function $f$ and $p_{i},i\in[t]$ are distinct primes.
If $k=1$, then the problem reduces to finding lower bound of Fourier coefficients on $\mathbb{Z}_{p}^{n}$, where $p$ is a prime number, which we already know to be equal to $\frac{1}{(p^{2}s_{f})^{\lceil(p-1)/2\rceil}}$ by Theorem 1.5.
Now, let us solve this for $k\geq 2$. We know by Theorem 2.1 that
$|\prod_{i_{1}=1}^{p_{1}-1}\cdots\prod_{i_{t}=1}^{p_{t}-1}g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})|\geq 1.$
Further, the conjugate of $g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})$, namely $\overline{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}$, is nothing but $=g(\omega_{p_{1}}^{p_{1}-i_{1}},\ldots,\omega_{p_{t}}^{p_{t}-i_{t}})$. Finally, since $|g|_{1}\leq p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}$, it follows that for any $i_{1}\in[p_{1}-1],\ldots,i_{t}\in[p_{t}-1]$,
[…]
Therefore,
$|\prod_{i_{1}=1}^{p_{1}-1}\cdots\prod_{i_{t}=1}^{p_{t}-1}g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})|\geq 1$
$\Rightarrow\prod_{i_{1}=1}^{(p_{1}-1)/2}\cdots\prod_{i_{t}=1}^{(p_{t}-1)/2}|g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})|^{2}\geq 1$
$\Rightarrow\prod_{i_{1}=1}^{(p_{1}-1)/2}\cdots\prod_{i_{t}=1}^{(p_{t}-1)/2}|\frac{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|^{2}\geq\biggl{(}\frac{1}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}\biggr{)}^{(p_{1}-1)\cdots(p_{t}-1)}$
$\Rightarrow|\frac{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|^{2}\geq\frac{1}{p_{1}^{m_{1}(p_{1}-1)}\cdots p_{t}^{m_{t}(p_{t}-1)}}$
$\Rightarrow|\frac{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|\geq\frac{1}{p_{1}^{m_{1}(p_{1}-1)/2}\cdots p_{t}^{m_{t}(p_{t}-1)/2}}$
$\Rightarrow|\frac{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|\geq\frac{1}{(p_{1}^{2}\cdots p_{t}^{2}s_{f})^{(p_{1}-1)\cdots(p_{t}-1)/2}}$
$\Rightarrow|\frac{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|\geq\frac{1}{(m^{2}s_{f})^{\varphi(m)/2}}\geq\frac{1}{(m^{2}s_{f})^{\lceil\varphi(m)/2\rceil}},$
where $m=p_{1}\cdots p_{t}$ and $\varphi(m)=(p_{1}-1)\cdots(p_{t}-1)$.
Combining the two cases, we have
$|\frac{g(\omega_{p_{1}}^{i_{1}},\ldots,\omega_{p_{t}}^{i_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|\geq\frac{1}{(m^{2}s_{f})^{\lceil\varphi(m)/2\rceil}}.$
Remark F.10.
Theorem F.8 implies that if a Fourier coefficient of a Boolean function f is $(m_{1},\ldots,m_{t})$ granular, where $m_{i}=\lceil\log_{p_{i}}s^{\frac{1}{t}}\rceil+1$ for all $i\in[t]$, then its absolute value is $\geq\frac{1}{(m^{2}s)^{\lceil\varphi(m)/2\rceil}}$. It directly implies that if a Fourier coefficient is $2(m_{1},\ldots,m_{t})$-granular, then its absolute value is $\geq\frac{1}{(m^{2}s)^{\varphi(m)}}$ (The case $k=1$ and $p_{1}=2$ needs to be taken care of separately, but we already did that in Remark 2.3).
Now let us prove Theorem F.4.
Proof F.11 (Proof of Theorem F.4).
Since we have already proved this theorem for $k=1$, that is for $\mathbb{Z}_{p}^{n}$ (and  [15] proved this for $\mathbb{Z}_{2}^{n}$), where $p$ is an odd prime number, we will prove for $k\geq 2$ here.
Let us assume $m_{i}=\lceil\log_{p_{i}}s^{\frac{1}{t}}\rceil+1$, for all $i\in[t]$. Let $B$ be the set of character functions corresponding to the $s$ largest coefficients and $S$ be the set of rest of the characters. Then, by Theorem F.2, for all $\chi_{r_{j}}\in B$, $\widehat{f}(\chi_{r_{j}})$ is $\frac{\mu}{\sqrt{s}}-$close to $(m_{1},\ldots,m_{t}))$-granular. Therefore, for all $\chi_{r_{j}}\in B$, we can write $f(\chi_{r_{j}})$ as the sum of $\widehat{F}(\chi_{r_{j}})$ and $\widehat{G}(\chi_{r_{j}})$, that is,
$\widehat{f}(\chi_{r_{j}})=\widehat{F}(\chi_{r_{j}})+\widehat{G}(\chi_{r_{j}}),$
where $|\widehat{F}(\chi_{r_{j}})|$ is $(m_{1},\ldots,m_{t}))$-granular and $|\widehat{G}(\chi_{r_{j}})|\leq\frac{\mu}{\sqrt{s}}$.
Now, for $\chi\in\mathcal{S}$, let us set $\widehat{F}(\chi)=0$ and $\widehat{G}(\chi)=\widehat{f}(\chi)$. Therefore, we can write $f(x)$ as sum of $F(x)$ and $G(x)$, that is,
$f(x)=F(x)+G(x).$
Clearly, $F$ is $s-$sparse and $|\widehat{F}(\chi)|$ is $(m_{1},\ldots,m_{t}))$-granular for all Fourier coefficients $\widehat{F}(\chi)$ of $F$.
Let us look at $G(x)$. We have,
$\mathbb{E}[G(x)^{2}]$ $=\mathbb{E}[\sum_{\chi\in\mathcal{B}}\widehat{G}(\chi)^{2}+\sum_{\chi\in\mathcal{S}}\widehat{G}(\chi)^{2}]$
$=\mathbb{E}[\sum_{\chi_{r_{j}}\in\mathcal{B}}\widehat{G}(\chi_{r_{i}})^{2}+\sum_{\chi\in\mathcal{S}}\widehat{f}(\chi)^{2}]$
$\leq s\cdot\frac{\mu^{2}}{s}+\mu^{2}$
$\leq 2\mu^{2}.$
Claim: Range of $F$ is {-1,+1}.
Proof F.12 (Proof of the claim).
Since $f$ is Boolean valued and $f=F+G$, so we have,
$1=f^{2}=(F+G)^{2}=F^{2}+2FG+G^{2}=F^{2}+G(2F+G)=F^{2}+G(2f-G).$ (F.7)
Let $H=G(2f-G)$. Let us look at the Fourier coefficients $\widehat{H}(\chi)$ of $H$.
$|\widehat{H}(\chi)|$ $\leq||G||_{2}||2f-G||_{2}$ by Lemma A.12
$\leq||G||_{2}(||2f||_{2}+||G||_{2})$
$\leq\sqrt{2}\mu\cdot 2\cdot 1+2\mu^{2}$
$<4\mu$
$\leq\frac{1}{2(m^{2}s)^{\varphi(m)}},$ (F.8)
where $m=p_{1}\cdots p_{t}$ and $\varphi(m)=(p_{1}-1)\cdots(p_{t}-1)$.
Now, since $\widehat{F}(\chi)$ is $(m_{1},\ldots,m_{t}))$-granular, any Fourier coefficient $\widehat{F^{2}}(\chi)$ of $F^{2}$ is $2(m_{1},\ldots,m_{t}))$-granular. That is, $\widehat{F^{2}}(\chi)=\frac{g^{\prime}(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{2m_{1}}\cdots p_{t}^{2m_{t}}}$, where $g^{\prime}(X_{1},\ldots,X_{t})\in\mathbb{Z}[X_{1},\ldots,X_{t}]$ is a nonzero multivariate polynomial with degree $\leq\varphi(m)$. Also, by Remark F.10, $|\widehat{F^{2}}(\chi)|\geq\frac{1}{(m^{2}s)^{\varphi(m)}}$.
Now, from Equation F.7, we have,
$\widehat{F^{2}}(\chi_{0})+\widehat{H}(\chi_{0})=1$
and
$\widehat{F^{2}}(\chi)+\widehat{H}(\chi)=0,\text{ for all }\chi\neq\chi_{0},$
where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathcal{G}$, as defined in Lemma A.2, part 1.
Since $\widehat{F^{2}}(\chi)\geq\frac{1}{(m^{2}s)^{\varphi(m)}}$ and $\widehat{H}(\chi)<\frac{1}{2(m^{2}s)^{\varphi(m)}}$, these two cannot sum up to 00, therefore, $\widehat{F^{2}}(\chi)=0$, $\widehat{H}(\chi)=0$ for all $\chi\neq 0$.
Now, let us look at $\widehat{F^{2}}(\chi_{0})$.
$\widehat{F^{2}}(\chi_{0})+\widehat{H}(\chi_{0})=1$ $\Rightarrow\widehat{H}(\chi_{0})=1-\frac{g^{\prime}(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{2m_{1}}\cdots p_{t}^{2m_{t}}}$
$\Rightarrow\frac{p_{1}^{2m_{1}}\cdots p_{t}^{2m_{t}}-g^{\prime}(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{2m_{1}}\cdots p_{t}^{2m_{t}}},$
which implies that $\widehat{H}(\chi_{0})$ is $2(m_{1},\ldots,m_{t})$-granular if $\widehat{H}(\chi_{0})\neq 0$. So, by Remark F.10, $|\widehat{H}(\chi_{0})|\geq\frac{1}{(m^{2}s)^{\varphi(m)}}$, which is not possible since $|\widehat{H}(\chi_{0})|\leq\frac{1}{2(m^{2}s)^{\varphi(m)}}$ by F.12. So, $\widehat{H}(\chi_{0})=0$, hence $\widehat{F^{2}}(\chi_{0})=1$. Therefore $F^{2}(x)=0$ for all $x\in\mathcal{G}$, which implies that the range of $F$ is $\{-1,+1\}$.
Now, let us find $\Pr_{x}[f(x)\neq F(x)]$. Since $F$ and $f$ both are Boolean valued, so the range of $G$ is equal to $\{-2,0,+2\}$. So,
$\Pr_{x}[f(x)\neq F(x)]$ $=\Pr_{x}[|G(x)|=2]$
$=\frac{1}{4}\mathbb{E}_{x}[G(x)^{2}]$ by Markov’s inequality
$\leq\frac{\mu^{2}}{2}.$
Therefore, by Lemma A.15, the $\ell_{2}$ distance between $f$ and $F$ is $\leq\sqrt{4\times\frac{\mu^{2}}{2}}=\sqrt{2}\mu$.
Appendix G Algorithm over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$
The algorithm for testing sparsity of Boolean-valued functions over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ is a simple generalization of the Algorithm 1. The main crucial difference is that the subspace $H$ defined in Line 5 has been defined by drawing a set of random subspaces $H_{1},\dots,H_{t}$ from $\mathbb{Z}_{p_{1}}^{n_{1}},\dots,\mathbb{Z}_{p_{t}}^{n_{t}}$ respectively, and finally defining $H=H_{1}\times\dots\times H_{t}$. The rest of the algorithm (except for slight changes to the parameters) remains the same.
The proof of correctness uses Theorem F.8. Just like Theorem 1.5 was the central technical result used in the proof of correctness of Algorithm 1, the Theorem F.8 performs the exact same role in the proof of correctness of the algorithm for Boolean-valued functions over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$.
We present the algorithm for Boolean-valued functions over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ in Algorithm 2. We skip the proof of the correctness of the algorithm.
Since the algorithm is already known for functions on $\mathbb{Z}_{2}^{n}$ (proved in  [15]) and $\mathbb{Z}_{p}^{n}$, where $p$ is a prime number (proved in Section E.1), so let us assume that $k\geq 2$.
Algorithm 2 General Test for Sparsity
1:  Input: $s,\epsilon$, and query access to $f:\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}\to\{-1,+1\}$.
2:  Output: YES, if $f$ is $s$-sparse, and NO, if it is $\epsilon$-far.
3:  Let $\mathcal{G}=\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ and $m=p_{1}p_{2}\dots p_{t}$
4:  Parameter setting:
5:  $t_{i}:=\lceil 2\log_{p_{i}}s^{1/t}+\log_{p_{i}}20\rceil+1$
6:  $\tau:=\min(\frac{\epsilon^{2}}{40\prod_{i}p_{i}^{t_{i}}},\frac{1}{(m^{2}s)^{\phi(m)}})$
7:  $M:=O(\log(\prod_{i}p_{i}^{t_{i}})\cdot 1/\tau^{2})$
8:  for $i=1$ to $t$ do
9:     Choose $v_{i,1},\dots,v_{i,j_{i}}$ linearly independent vectors uniformly at random from $\mathbb{Z}_{p_{i}}^{n_{i}}$.
10:     Let $H_{i}=\operatorname{Span}\{v_{i,1},\dots,v_{i,j_{i}}\}$
11:  end for
12:  $H=H_{1}\times\dots\times H_{t}$
13:  Pick $(z_{1},x_{1}),\dots,(z_{M},x_{M})$ uniformly and independently from $H\times\mathcal{G}$
14:  Query $f(x_{1}),\dots,f(x_{M})$ and $f(x_{1}-z_{1}),\dots,f(x_{M}-z_{M})$
15:  for For every $r\in\mathcal{G}$ do
16:     Let $\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$ be the estimate of $wt(r+H^{\perp})$
17:  end for
18:  if number of $r$ for which the estimate of $wt(r+H^{\perp})$ is $\geq 2\tau/3$ is $\leq s$ then
19:     Output YES
20:  else
21:     Output NO
22:  end if
Appendix H Degree over $\mathbb{Z}_{p}$
Definition H.1.
Let $r_{1},\ldots,r_{k}\in\mathbb{Z}_{p}^{n}$ and $b=(b_{1},\ldots,b_{t})\in\mathbb{Z}_{p}^{t}$. We define the subspace $V_{b,r_{1},\ldots,r_{k}}$ by
$V_{b,r_{1},\ldots,r_{k}}=\{x\in\mathbb{Z}_{p}^{n}:r_{j}\cdot x=b_{j}(\mod p)\ \forall j\in[t]\}.$ (H.1)
The following definition is a generalization of $\deg_{2}(f)$, that is degree over $\mathbb{Z}_{2}$ of a Boolean function $f$, as defined in [7].
Definition H.2.
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean valued function. Consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a subspace of $\mathbb{Z}_{p}^{n}$ as defined in Definition H.1. Then the degree over $\mathbb{Z}_{p}$ of $f$, denoted by $\deg_{p}$, is defined in the following way.
$\deg_{p}(f)=\max_{\ell}\{\ell=\dim(V_{b,r_{1},\ldots,r_{t}}):s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}=p^{\dim(V_{b,r_{1},\ldots,r_{t}})}\},$
where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$.
Remark H.3.
In Definition H.2, $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}=p^{\dim(V_{b,r_{1},\ldots,r_{t}})}$ means that the restriction $f|_{V_{b,r_{1},\ldots,r_{t}}}$ has full sparsity.
We will often refer to the dimension of a Boolean valued function $f$ as $\dim(f)$, which is equal to $\dim(\operatorname{supp}(\widehat{f}))$.
Next, let us define the $p\times p$ matrix $V_{1}$ by
$V_{1}=\begin{bmatrix}1&1&1&\cdots&1\\ 1&\omega_{p}&\omega_{p}^{2}&\cdots&\omega_{p}^{p-1}\\ 1&\omega_{p}^{2}&\omega_{p}^{4}&\cdots&\omega_{p}^{2(p-1)}\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ 1&\omega_{p}^{p-1}&\omega_{p}^{2(p-1)}&\cdots&\omega_{p}^{(p-1)(p-1)}\\ \end{bmatrix}$
and inductively define the $p^{n}\times p^{n}$ matrices $V_{n}$ by
$V_{n}=\begin{bmatrix}V_{n-1}&V_{n-1}&V_{n-1}&\cdots&V_{n-1}\\ V_{n-1}&\omega_{p}V_{n-1}&\omega_{p}^{2}V_{n-1}&\cdots&\omega_{p}^{p-1}V_{n-1}\\ V_{n-1}&\omega_{p}^{2}V_{n-1}&\omega_{p}^{4}V_{n-1}&\cdots&\omega_{p}^{2(p-1)}V_{n-1}\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ V_{n-1}&\omega_{p}^{p-1}V_{n-1}&\omega_{p}^{2(p-1)}V_{n-1}&\cdots&\omega_{p}^{(p-1)(p-1)}V_{n-1}\\ \end{bmatrix}$
for all positive integer $n$.
Alternately, one can define $V_{n}$ by taking $n$ many Kronecker products of $V_{1}$, i.e., $V_{n}:=V_{1}^{\otimes n}=V_{1}\otimes\cdots\otimes V_{1}$.
Lemma H.4.
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean valued function and $\widehat{f}$ denotes its Fourier transform. Then $\widehat{f}=\frac{1}{p^{n}}V_{n}f$, where $f$ is the $p^{n}\times 1$ matrix formed by the values $f(x)$, $x\in\mathbb{Z}_{p}^{n}$ are arranged in usual lexicographic order, and $\widehat{f}$ is the $p^{n}\times 1$ matrix formed by the values $\widehat{f}(\chi_{r})$, $r\in\mathbb{Z}_{p}^{n}$ are arranged in usual lexicographic order.
Proof H.5.
We will use induction on $n$ to prove this.
Base case.
When $n=1$, then for each element $i\in\mathbb{Z}_{p}$, the Fourier coefficients are given by
$\widehat{f}(\chi_{i})=\frac{1}{p}\biggl{(}f(0)+f(1)\omega_{p}^{i}+f(2)\omega_{p}^{2i}+\ldots+f(p-1)\omega_{p}^{(p-1)i}\biggr{)}.$
Hence $\widehat{f}=\frac{1}{p}V_{1}f$ in this case.
Induction step.
Let for all $n\leq m-1$, $\widehat{f}=\frac{1}{p^{n}}V_{n}f$.
For $n=m$, for each element $(i,r)\in\mathbb{Z}_{p}\times\mathbb{Z}_{p}^{m-1}$, the Fourier coefficients of $f$ are given by
$\widehat{f}(\chi_{(i,r)})=\frac{1}{p^{m}}\sum_{x\in\mathbb{Z}_{p}^{m-1}}\biggl{(}\bigl{(}f(0,x)+f(1,x)\omega_{p}^{i}+f(2,x)\omega_{p}^{2i}+\cdots+f(p-1,x)\omega_{p}^{(p-1)i}\bigr{)}\omega_{p}^{r\cdot x}\biggr{)}.$ (H.2)
From Equation H.2, any block corresponding to the coefficients $\widehat{f}(\chi_{i,r})$ and the function values $f(j,x)$ is given by the matrix $\omega_{p}^{ij}V_{m-1}$, $r,x\in\mathbb{Z}_{p}^{m-1}$. Hence, $\widehat{f}=\frac{1}{p^{m}}V_{m}f$.
Thus by induction it follows that $\widehat{f}=\frac{1}{p^{n}}V_{n}f$ for any positive integer $n$.
One can equivalently define $\deg_{p}$ in the following way.
Definition H.6.
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean valued function. Let us arrange the elements $x\in\mathcal{G}$ in lexicogarphic order along the rows and columns of the matrix $V_{n}$. Consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a subspace of $\mathbb{Z}_{p}^{n}$ as defined in Definition H.1. Then the degree over $\mathbb{Z}_{p}$ of $f$ is defined as the dimension of the largest subspace $V_{b,r_{1},\ldots,r_{t}}$ such that the restriction $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value $1$ at those points such that each entry of the row/column matrix formed by the sum of the rows/columns corresponding to the elements where $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value $1$ is $\neq 0$.
[Note that in that case all the Fourier coefficients are nonzero.]
Remark H.7.
Observe that any character of $f$ can be written as $\frac{1}{p^{n}}\biggl{(}\sum_{i}c_{i}\omega_{p}^{i}-\sum_{j}d_{j}\omega_{p}^{j}\biggr{)}$, where $c_{i},d_{j}$ are non-negative integers for all $i,j$ and $\omega_{p}$ is a primitive root of unity. Definition H.6 follows from the following two facts:
  • $a_{0}+a_{1}\omega_{p}+a_{2}\omega_{p}^{2}+\cdots+a_{p-1}\omega_{p}^{p-1}=0$ if and only if $a_{0}=a_{1}=a_{2}=\cdots=a_{p-1}$, when $p$ is a prime number.
  • Each root of unity appears $p^{n-1}$ number of times in each rows/columns of the matrix $V_{n}$, which implies that the sum of elements of each row/column is equal to 00.
These two facts together imply that, for a column with $\sum_{i}c_{i}\omega_{p}^{i}\neq 0$, we have the corresponding character $=\frac{1}{p^{n}}\biggl{(}\sum_{i}c_{i}\omega_{p}^{i}-\sum_{j}d_{j}\omega_{p}^{j}\biggr{)}=\frac{1}{p^{n}}\biggl{(}\sum_{i}c_{i}\omega_{p}^{i}+\sum_{i}c_{i}\omega_{p}^{i}\biggr{)}\neq 0$.
Appendix I Lower bound
In this section, we will prove Theorem 1.9, that is, find a lower bound on the number of queries for our testing algorithm.
Theorem I.1 (Theorem 1.9, restated).
For adaptively testing $s$-sparsity, the query lower bound of any algorithm is $\Omega(\sqrt{s})$.
Let $\tau>0$ and $C$ be a constant dependent on $\tau$, and will be determined later. Let us choose a random $t$-dimensional coset structure $(H,\mathcal{C})$ in $\mathbb{Z}_{p}^{Ct}$. That is, $H$ is a random subspace of $\mathbb{Z}_{p}^{Ct}$ of dimension $t$, and $\mathcal{C}$ is the set of all cosets of $H$. Now, let us define the following two probability distributions:
  • Let us construct random functions $f$ by making $f$ constant on each coset of $\mathcal{C}$, the constant being chosen randomly from $\{-1,+1\}$. This implies that the dimension of $f$ is $t$. Let us call this probability distribution $\mathcal{D}_{yes}$.
  • We choose a random function $f$ randomly from $\mathbb{Z}_{p}^{Ct}$, conditioned on the fact that $f$ is $2-\tau$ far in $\ell_{2}$ from any function which has $\deg_{p}=t$. Let us call this distribution $\mathcal{D}_{No}$.
We will prove the following theorem first, Theorem 1.9 follows as a corollary of this.
Theorem I.2.
Any adaptive query algorithm that distinguishes $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ with probability $\geq\frac{1}{3}$ has to make at least $\Omega(p^{\frac{t}{2}})$ queries.
The following lemma will be useful in proving Theorem I.2.
Lemma I.3.
For any Boolean valued function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$,
$p^{\dim(f)}\geq s_{f}\geq p^{\deg_{p}(f)}.$
Proof I.4.
The proof of the first inequality follows trivially from the definitions of sparsity $s_{f}$ (Definition A.9) and dimension $\dim(f)$ (Remark H.3) of $f$.
The proof of the second inequality follows from the definition of $\deg_{p}(f)$ (Definition H.2), since $s_{f}\geq s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}=p^{\dim(V_{b})}=p^{\deg_{p}(f)}$, where $V_{b}$ is the largest subspace such that $f|_{V_{b,r_{1},\ldots,r_{t}}}$ has full sparsity.
Corollary I.5 (Corollary of Lemma I.3).
The set of functions with $\dim=t$ is a subset of the set of functions with sparsity $s_{f}=p^{t}$, which in turn is a subset of the set of functions with $\deg_{p}=t$.
Theorem 1.9 directly follows from Theorem I.2 and Lemma I.3.
We will show that if an adaptive query algorithm makes less than $q<\Omega(p^{t/2})$ queries, then the total variation distance $||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}$ between the two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ is $\leq\frac{1}{3}$, where the total variation distance is as defined in Definition A.17 and Lemma A.18. This proves that any adaptive query algorithm which distinguishes between $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$, that is, for which $||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}>\frac{1}{3}$ must make at least $\Omega(p^{t/2})$ queries.
Any adaptive query algorithm that makes $q$ queries can be thought of as a decision tree, whose internal nodes are elements from $\mathbb{Z}_{p}^{Ct}$ and the leaves are labeled by accept or reject. Also, the length of each path from the root to the leaves of the tree is equal to $q$ and no element repeats in the tree, that is, it does not appear more than once in the tree.
Let $\mathcal{L}_{Yes}$ and $\mathcal{L}_{No}$ be the distributions on the leaves of the decision tree from $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ respectively. We will show that $||\mathcal{L}_{Yes}-\mathcal{L}_{No}||_{TV}\leq\frac{1}{3}$ when the number of queries $q\leq\frac{1}{\sqrt{20}}p^{\frac{t}{2}}$.
Lemma I.6.
$||\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||\leq\frac{1}{19}$.
Proof I.7 (Proof of Lemma I.6).
Consider the set of all functions of dimension $t$. Let the depth $q$ of the decision tree corresponding to our algorithm is $\leq\frac{1}{\sqrt{20}}p^{\frac{t}{2}}$. Then $t\geq 2\log_{p}q+\log_{p}20$. So by Theorem E.13 (part 2), all the characters on a particular path fall into different buckets with probability $\geq(1-\frac{1}{20})$. Let us choose a function $f$ randomly from $\mathcal{D}_{Yes}$. For each leaf $\ell$, let us consider the characters on the path from root to $\ell$. Since the length of the path is $q$, so with probability $1-\frac{1}{20}$, the characters on the path belongs to different buckets. Then the probability that $f$ is consistent with the path to $\ell$, given that all the characters on the path from root to $\ell$ belong to different buckets is $=\frac{1}{p^{q}}$. Therefore, the probability that $f$ is consistent with the path to $\ell$ is $\geq(1-\frac{1}{20})\frac{1}{p^{q}}=\frac{19}{20}\frac{1}{p^{q}}$. Let us denote this probability by $\mathcal{P}_{Yes}$.
Now, for each path from root to $\ell$, we know that $\mathcal{P}_{Uniform}(x)=\frac{1}{p^{q}}$, where $x$ is a character in the path from root to $\ell$. Then,
$||\mathcal{P}_{Yes}-\frac{19}{20}\mathcal{P}_{Uniform}||_{1}$
=$=\sum_{\{x\in\text{ the path to }\ell\}}|\mathcal{P}_{Yes}(x)-\frac{19}{20}\mathcal{P}_{Uniform}(x)|$
=$=\sum_{\{x\in\text{ the path to }\ell\}}|\mathcal{P}_{Yes}(x)-\frac{19}{20}\frac{1}{p^{q}}|$
=$=\sum_{\{x\in\text{ the path to }\ell:\mathcal{P}_{Yes}(x)\geq\frac{19}{20}\frac{1}{p^{q}}\}}|\mathcal{P}_{Yes}(x)-\frac{19}{20}\frac{1}{p^{q}}|$ since $\text{since }\mathcal{P}_{Yes}(x)\geq\frac{19}{20}\frac{1}{p^{q}}$
[$\leq\biggl{[}\sum_{\{x\in\text{ the path to }\ell:\mathcal{P}_{Yes}(x)\geq\frac{19}{20}\frac{1}{p^{q}}\}}\mathcal{P}_{Yes}(x)\biggr{]}-\frac{19}{20}$
$\leq 1-\frac{19}{20}=\frac{1}{20}.$
Therefore,
$||\mathcal{P}_{Yes}-\mathcal{P}_{Uniform}||_{1}$ $\leq||\mathcal{P}_{Yes}-\frac{\mathcal{P}_{Yes}}{\frac{19}{20}}||_{1}+||\frac{\mathcal{P}_{Yes}}{\frac{19}{20}}-\mathcal{P}_{Uniform}||_{1}$
$\leq\frac{1}{19}||\mathcal{P}_{Yes}||_{1}+\frac{\frac{1}{20}}{\frac{19}{20}}$
$=\frac{2}{19}$ since $\text{since }||\mathcal{P}_{Yes}||_{1}=1.$
Hence,
$||\mathcal{P}_{Yes}-\mathcal{P}_{Uniform}||_{TV}=\frac{1}{2}||\mathcal{P}_{Yes}-\mathcal{P}_{Uniform}||_{1}=\frac{1}{19}.$
Lemma I.8.
Let $\mathcal{D}_{Uniform}$ be the uniform distribution on the set of functions from $\mathbb{Z}_{p}^{Ct}$ to $\{-1,+1\}$. Then, the probability that the degree over $\mathbb{Z}_{p}$ of a randomly chosen function $f$ is $(2-\tau)$ close to $t$ in $\ell_{2}$ is $\leq\frac{1}{10}$.
Proof I.9 (Proof of Lemma I.8).
Let us fix a function $g$ from $\mathbb{Z}_{p}^{Ct}$ to $\{-1,+1\}$ such that $\deg_{p}(g)=t$. Let $f$ be randomly chosen function, where $f$ is also from $\mathbb{Z}_{p}^{Ct}$ to $\{-1,+1\}$. We need to show that the probability that the $\ell_{2}$ distance between $f$ and $g$ is $\leq$ $(2-\tau)$ is $\leq\frac{1}{10}$.
Let $\mathbb{I}_{x}$ be the indicator variable that takes the value $1$ when $f(x)\neq g(x)$, and it is 00 otherwise. Then $X_{g}^{\prime}=\frac{\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{I}_{x}}{p^{Ct}}$ denotes the Hamming distance between them. Also, by Lemma A.15, if $X_{g}$ denotes thesquare of the $\ell_{2}$ distance between $f$ and $g$, then $X_{g}=4X_{g}^{\prime}$. Now,
$\mathbb{E}[X_{g}^{\prime}]=\frac{\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{E}[\mathbb{I}_{x}]}{p^{Ct}}=\frac{1}{p^{Ct}}\frac{1}{2}p^{Ct}=\frac{1}{2}$
$\Rightarrow\mathbb{E}[X_{g}]=4\mathbb{E}[X_{g}^{\prime}]=2.$
Therefore, by Chernoff (Lemma A.21),
$\Pr[X_{g}^{\prime}\leq\frac{1}{2}-\frac{\tau}{4}]$ $\leq\Pr[|X_{g}^{\prime}-\mathbb{E}[X_{g}^{\prime}]|\geq\frac{\tau}{4}]$
$=\Pr[|\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{I}_{x}-\mathbb{E}[\sum_{x\in\mathbb{Z}_{p}^{Ct}}\mathbb{I}_{x}]|\geq\frac{\tau}{4}\times p^{Ct}]$
$\leq 2\exp(-\frac{2}{3}\bigl{(}\frac{\tau}{4}\bigr{)}^{2}p^{Ct}),$
which implies that
$\Pr[X_{g}\leq 2-\tau]\leq 2\exp(-\frac{1}{24}\tau^{2}p^{Ct}).$
Now, the number of functions with degree over $\mathbb{Z}_{p}$ equal to $t$ is $\leq p^{\binom{Ct}{t}}$. The probability that $f$ is $(2-\tau)$-close to a degree $t$ function in $\ell_{2}$ is given by,
$\Pr[\cup_{g:\deg(g)=t}\{X_{g}\leq 2-\tau\}]$
$\leq p^{\binom{Ct}{t}}\times 2\exp(-\frac{2}{3}\tau^{2}p^{Ct})$
$\leq 2\exp\biggl{(}(aC)^{t}\ln p-\frac{2}{3}\tau^{2}p^{Ct}\biggr{)}$ by Stirling’s approximation
$\leq\frac{1}{10},$
taking $C\geq\frac{1}{t\ln p}\ln\biggl{(}\frac{3((aC)^{t}\ln p+\ln 20)}{2\tau^{2}}\biggr{)}$, where $a$ is a constant.
Corollary I.10.
$||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}$.
Proof I.11.
Let $A$ be the set of all functions $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ which are $(2-\tau)$ close to $\mathbb{Z}_{p}$-degree $t$ in $\ell_{2}$. Then $\mathcal{P}_{No}(A)=0$, and by Lemma I.8, $\mathcal{P}_{Uniform(A)}\leq\frac{1}{10}$, where $\mathcal{P}_{No}$ and $\mathcal{P}_{Uniform}$ denote the probability over the distributions $\mathcal{D}_{No}$ and $\mathcal{D}_{Uniform}$ respectively. So, $|\mathcal{L}_{No}(A)-\mathcal{L}_{Uniform}(A)|\leq\frac{1}{10}$.
Hence,
$||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}.$
Proof I.12 (Proof of Theorem I.2).
From Lemma I.6, we have $||\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||\leq\frac{1}{19}$. Also, from Corollary I.10, we have $||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}\leq\frac{1}{10}$. Therefore, by triangle inequality,
$||\mathcal{L}_{Yes}-\mathcal{L}_{No}||_{TV}$
$\leq||\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||_{TV}+||\mathcal{L}_{No}-\mathcal{L}_{Uniform}||_{TV}$
$\leq\frac{1}{19}+\frac{1}{10}\leq\frac{1}{3}.$
Hence,
$||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}=||\mathcal{L}_{Yes}-\mathcal{L}_{No}||_{TV}\leq\frac{1}{3}.$
Remark I.13.
The proof of Theorem I.2 (hence Theorem 1.9) holds for $\mathcal{G}$, since $\deg_{p}$ of a Boolean valued function over $\mathbb{Z}_{p}^{n}$ can be defined for a Boolean valued function over $\mathcal{G}$ in a similar manner, and the proof goes along the lines of the proof of Theorem I.2.